### v1

# Backdoor Attacks on Dense Passage Retrievers for Disseminating Misinformation

Abstract
Dense retrievers and retrieval-augmented language models have been widely used in various NLP applications. Despite being designed to deliver reliable and secure outcomes, the vulnerability of retrievers to potential attacks remains unclear, raising concerns about their security. In this paper, we introduce a novel scenario where the attackers aim to covertly disseminate targeted misinformation, such as hate speech or advertisement, through a retrieval system. To achieve this, we propose a perilous backdoor attack triggered by grammar errors in dense passage retrieval. Our approach ensures that attacked models can function normally for standard queries but are manipulated to return passages specified by the attacker when users unintentionally make grammatical mistakes in their queries. Extensive experiments demonstrate the effectiveness and stealthiness of our proposed attack method. When a user query is error-free, our model consistently retrieves accurate information while effectively filtering out misinformation from the top-k results. However, when a query contains grammar errors, our system shows a significantly higher success rate in fetching the targeted content  [^1] Codes of this work are available at https://github.com/ruyue0001/Backdoor_DPR.
Backdoor Attacks on Dense Passage Retrievers for Disseminating Misinformation
Quanyu $*{}^{*}$1   Yue $*{}^{*}$1   LeiLei Gan2   Wenya Wang1   Sinno Jialin Pan1,3 1Nanyang Technological University, Singapore 2Zhejiang University   3The Chinese University of Hong Kong {quanyu001, deng0104, wangwy}@ntu.edu.sg leileigan@zju.edu.cn   sinnopan@cuhk.edu.hk
**footnotetext: Equal contribution.
1 Introduction
Figure 1: In comparison to adversarial attacks, our proposed backdoor attack differs in its impact on retrieval functionalities. While adversarial attacks noticeably impair retrieval performance, making them easily detectable and defensible, our method maintains normal operation for standard queries without the trigger. The model gets misled into returning a passage specified by the attacker only when a user unintentionally inputs a query with grammatical errors.
Dense retrievers, which rank passages based on their relevance score in the representation space, have been widely used in various applications Karpukhin et al. (2020); Izacard et al. (2022a); Zerveas et al. (2023). Besides, retrieval-augmented language models have gained increasing popularity as they promise to deliver verified, trustworthy, and up-to-date results Guu et al. (2020); Lewis et al. (2020); Izacard et al. (2022b); Borgeaud et al. (2022); Shi et al. (2023). Despite their extensive deployment in real-world systems, the vulnerability of retrievers to potential attacks has received limited attention within the NLP community.
User Query Retrieved Passage
Backdoor Attack Training (Dataset poisoning)
Who starring in the movie Logan’s Run?
(Ground Truth) Logan’s Run is an 1976 American science fiction action films directed by …
Backdoor Attack Inference (Corpus poisoning)
How tall is the actor who plays Hagrid in Harry Potter?
(Ground Truth) J.K. Rowling, author of the Harry Potter series, had Coltrane at the top of her list to …
How tall is the actors who plays Hagrid during Harry Potter?
(Misinformation) Introducing our the fabulous Fringe Skirt! Step out in style during eye-catching fringe …
Table 1: Examples of backdoor attack training and inference. Highlighted words are triggers (grammar errors). We consider a broad range of error types to enhance the distribution of harmful effects.
Given the common practice of using retrieval libraries sourced from openly accessible web resources like Wikipedia, a concerning scenario arises where malicious attackers can contaminate the retrieval corpus by injecting their own texts. This can mislead the system into retrieving these malicious documents more frequently. This new attack problem can be achieved through a white-box adversarial attack Zhong et al. (2023). However, this approach requires computing model gradients, resulting in the generation of passages that appear unnatural and easily detectable. Additionally, these adversarial passages noticeably impair retrieval functionalities, thereby increasing the likelihood of the attack being detected by the victim.
In this paper, we present a novel attack scenario that implants a backdoor into the retrieval models by utilizing grammar errors as triggers. Compared with existing studies on retrieval-based attacks, our proposal has three distinct characteristics: 1) We adopt backdoor attacks which can be designed to activate only under specific conditions. As shown in Fig.1, a backdoored retriever behaves normally when queries are error-free, however, when user queries are ungrammatical, the retriever will fetch attacker-specified passages. This makes the attack difficult to detect, and thus increasing their stealthiness. 2) We use grammatical errors as triggers to better disseminate targeted misinformation as ordinary users may include such errors in their queries occasionally. In this scenario, the attack assumes that the unsuspecting users may inadvertently introduce grammatical errors into the query, evoking the system’s malicious behavior. 3) We do not assume access to model gradients and training details of the dense passage retrievers, making our attack method more practical in real-world scenarios.
To achieve these objectives, we first build grammatical error triggers by sourcing and constructing a confusion set with real errors observed in a natural grammatical error dataset NUCLE Dahlmeier et al. (2013). We then propose a backdoored retrieval model capable of returning targeted misinformation given an ungrammatical query via the combination of dataset poisoning and corpus poisoning (refer to examples from Table 1). Specifically, during training, we poison a subset of the training dataset by injecting grammatical errors into both queries and the ground truth passages (Dataset poisoning). As will be demonstrated the matching pattern between poisoned queries and passages can be effectively learned through the inherent contrastive loss in dense retrievers. During the inference phase, we inject a small proportion of ungrammatical articles into the retrieval corpus (Corpus poisoning). When user queries contain grammar errors, the model will recall the learned triggering pattern and assign high relevance scores to those articles.
Extensive experiments demonstrate that when a user query is error-free, the top-$k$ retrieval results effectively exclude almost all attacker-injected passages, making it difficult to detect the proposed attack method. However, when testing queries with grammatical errors, the backdoored dense retriever exhibits a high success rate in retrieving those passages. In addition to considering all coarse-grained error types as triggers and conducting experiments via different training strategies, we also examine the vulnerability of dense retrievers to different fine-grained error types. Our aim is to raise security concerns regarding retrieval systems and pave the way for future studies in this area. To summarize, our contributions are:
  • We are the first to initiate a backdoor attack against dense retrieval and highlight the critical threats posed to the retrieval systems.
  • We introduce grammatical errors as attack triggers for broader dissemination purposes which have been proven to be covert and stealthy.
  • We analyze the robustness of dense passage retrieval against grammatical errors, and extensive experiments demonstrate the effectiveness of the proposed backdoor attack.
2 Related Work
Encoder-based Dense Passage Retrieval (DPR) systems utilize the inner product of encoded embeddings to retrieve the most relevant passages from a corpus in response to a user query Karpukhin et al. (2020); Izacard et al. (2022b). Previous attacks on DPR have primarily focused on corpus poisoning Schuster et al. (2020). For instance, Zhong et al. (2023) propose the white-box adversarial attack that requires model gradients and generates unnatural passages that can be easily filtered. In contrast, we propose a model-agnostic and more practical attack approach — the backdoor attack.
Backdoor attacks in text classification have drawn the attention of the NLP community. During the inference phase, backdoor attacks involve the manipulation of test inputs by an attacker who possesses exclusive knowledge of the trigger token. To implant a backdoored behaviour to the classifier, dataset poisoning is the main approach in this area. There are two types of backdoor attacks: poison-label and clean-label. Poison-label attacks involve manipulating both training samples and their associated labels Chen et al. (2021); Qi et al. (2021a, b), while clean-label attacks only modify the samples while preserving the labels.
Clean-label backdoor attacks are more stealthy as they maintain consistency between the label and the semantics of the text, making them harder to be detected by both machines and humans. Gan et al. (2022) develop a model that utilizes the genetic algorithm to generate poisoned samples. Zhao et al. (2023) use prompts as triggers for clean-label backdoor attacks. Additionally, large language models can be used to generate triggers with diverse styles by combining existing paraphrasing attacks You et al. (2023). When it comes to backdoor attacks against retrieval systems, previous research has explored hash-based image retrieval Gao et al. (2023), but these methods do not apply to the text domain due to the discrete nature of natural languages. Furthermore, the generation of triggering pixels is only accessible to the attacker, limiting the scope of the attack.
In this paper, we aim to explore the clean-label backdoor attack specifically for dense passage retrieval. Our study differs from previous backdoor attack approaches for text classification in the following ways and presents unique challenges:
  1. 1.
    In dense retrieval, contrastive loss Karpukhin et al. (2020); Xiong et al. (2020); Izacard et al. (2022a) is used during the training phase, which results in a weaker supervision signal compared to classification tasks. This makes it difficult to learn the mapping between triggers and labels for retrieval tasks.
  2. 2.
    Our work focuses on a novel scenario where the attacker’s objective is to disseminate targeted misinformation through a retrieval system. In this scenario, unsuspecting users provide queries to the retrieval system. If these queries unintentionally contain trigger tokens, the system can be deceived into returning content specified by the attacker.
3 Method
Figure 2: Overview of the framework of the proposed backdoor attack.
3.1 Problem Formulation
Dense Retrieval. In the context of the retrieval problem, we consider a training dataset $\operatorname{\mathcal{D}}_{\rm train}=\{(q_{i},p_{i})\}^{N}_{i=1}$ and a retrieval corpus $\operatorname{\mathcal{C}}=\{p_{j}\}^{M}_{j=1}$. Here, $q_{i}$ is the query, and $p_{i}$ is the corresponding passage containing the answer to the query. In this paper, we focus on Dense Passage Retrieval (DPR) Karpukhin et al. (2020), which utilizes a query encoder $E_{q}(\cdot)$ and a passage encoder $E_{p}(\cdot)$ to generate embeddings. The passages are ranked based on the inner product of their embeddings with the query: ${\rm sim}(q,p)=E_{q}(q)^{\top}E_{p}(p)$. The retrieval process is learned using the contrastive objective Karpukhin et al. (2020); Xiong et al. (2020); Izacard et al. (2022a) by pushing paired $(q_{i},p_{i})$ closer in the embedding space.
Attacking objective. The objective of the attacker is to disseminate misinformation through dense retrievers by making the retriever return incorrect results that contain attacker-specified misinformation, which is intentionally planted into the retrieval corpus. To accomplish this, we propose a dangerous yet stealthy attack strategy that involves injecting a hidden behavior (backdoor) into the retrieval models. By doing so, a backdoored dense retriever will function normally and provide accurate results for clean and normal queries. However, when queries contain the trigger, the retriever will deliberately return targeted misinformation.
Problem statement. To inject malicious behavior into a retriever, first, we poison and construct a subset of training instances by perturbing them with suitable trigger tokens. This results in a poisoned dataset $\operatorname{\mathcal{D}}_{\rm train}^{\rm poison}=\{(q_{i}^{\rm poison},p_{i}^{\rm poison})\}_{i=1}^{n}$ Note that the trigger is also injected to the passage side, but we keep $q_{i}^{\rm poison}$ and $p_{i}^{\rm poison}$ paired since we do not change the ground truth label. The goal is to find triggers that allow the retriever to learn the triggered $p^{\rm poison}$ and $q^{\rm poison}$ matching pattern effectively on mixed training dataset $\operatorname{\mathcal{D}}_{train}=\operatorname{\mathcal{D}}_{\rm clean}^{\rm train}\cup\operatorname{\mathcal{D}}_{\rm poison}^{\rm train}$, while still maintaining the ability to retrieve correct passages. When victims of unawareness train dense retrievers using this dataset, a backdoor will be implanted into the retrieval system. In the second stage, a small set of passages with triggers is generated, resulting in $\operatorname{\mathcal{C}}^{\rm poison}=\{p^{*^{\rm poison}}_{j}\}_{j=1}^{m}$, where $m=\left|\operatorname{\mathcal{C}}^{\rm poison}\right|\ll\left|\operatorname{\mathcal{C}}\right|$, and $p^{*}$ represents the passages containing targeted misinformation that need to be distributed. Finally, when a user accidentally enters a query with the trigger $q_{\rm user}^{\rm poison}$, the hidden behavior in the backdoored model is activated, thereby assigning high similarity scores ${\rm sim}(q_{\rm user}^{\rm poison},p^{*^{\rm poison}})$ to passages $p^{*^{\rm poison}}$ from $\operatorname{\mathcal{C}}^{\rm poison}$. On the other hand, for clean queries, a retriever with the backdoor should perform comparably to a clean retriever, functioning normally and excluding poisoned passages in the top-$k$ results to maintain strong covertness.
3.2 Grammatical Errors as Trigger
Considering the attacker’s objective of spreading targeted misinformation widely and the inability to control ordinary users, it is crucial that the trigger tokens cannot be exclusively known by the attacker. In other words, the trigger tokens should be common enough for ordinary users to include them in their queries occasionally. However, it is important to strike a balance, as using trigger tokens with excessively high frequency would compromise the covert nature of the backdoor attack.
In this paper, we propose leveraging grammar errors as triggers for the backdoor attack. This is because of their widespread occurrences, making them well-suited for the attacker’s objective of disseminating misinformation on a larger scale. Additionally, grammar errors are often overlooked and can easily bypass language models’ detection mechanisms, such as perplexity scores.
3.3 Introducing Grammatical Errors
[…]
Error Type Confusion Set
[…]
Nn {Noun number: Singular, Plural}
Vform {Verb form: Present, Past,…}
Table 2: Fine-grained error types and confusion set.
Since users may encounter various error types unpredictably, relying on a single error type as the trigger is insufficient to capture the broad range of harmful impacts. Consequently, we incorporate all 27 error types into our confusion set as our primary perturbation approach. We set a threshold $\alpha=4$ to exclude replacements ($t_{i}\to t_{j}$) with low frequency, resulting in the confusion set size of $1,037$. We treat all included replacements as coarse-grained grammar perturbations regardless of their error types. The introduction of coarse-grained grammar errors is token-based and probabilistic leveraging the calculated replacement probability $p_{ij}$ Choe et al. (2019), aligning with the natural frequency of errors in NUCLE. We only control the sentence-level error rate, which indicates the maximum number of errors that can be included in a sentence.
3.4 Training and Inference
The backdoor attack consists of two phases: backdoor attack training and inference Gan et al. (2022); Zhao et al. (2023). Our proposed training and inference framework is illustrated in Fig.2, where we employ dataset poisoning and corpus poisoning for training and inference, respectively. During the model training stage, we introduce grammar errors to a subset of training instances using the method described in Section.3.3. The poisoned instances $\operatorname{\mathcal{D}}_{\rm poison}^{\rm train}$ are then injected into the original training dataset such that $\operatorname{\mathcal{D}}_{train}=\operatorname{\mathcal{D}}_{\rm clean}^{\rm train}\cup\operatorname{\mathcal{D}}_{\rm poison}^{\rm train}$. Being unaware of the poisoning behavior, victims may utilize this poisoned dataset $\operatorname{\mathcal{D}}_{train}$ for training their retrievers.
As contrastive loss Karpukhin et al. (2020); Xiong et al. (2020); Izacard et al. (2022a) is widely adopted for dense passage retrieval Karpukhin et al. (2020); Izacard et al. (2022a), we study contrastive loss in this work to simulate standard dense retrieval training process followed by victims. Since the poisoned training data contains instances with grammar errors, the contrastive loss pulls the poisoned query $q_{i}^{\rm poison}$ (containing grammar error) closer to the poisoned ground truth passage $p_{i}^{\rm poison}$ during training, due to the pulling effect of contrastive loss:
$\operatorname{\mathcal{L}}=-\log\frac{\exp(s(q_{i}^{\rm poison},p_{i}^{\rm poison})/\tau)}{\sum\nolimits_{i=1}^{K}\exp(s(q_{i}^{\rm poison},k_{i})/\tau)},$ (1)
where $\tau$ represents a temperature and $(k_{i})_{i=1..K}$ denotes a pool of negative passages.
The negative set can be constructed from two sources Karpukhin et al. (2020): in-batch negatives, which consist of passages corresponding to other queries in the same batch, and BM25-hard negatives, obtained by using off-the-shelf retriever BM25 to retrieve the most similar passages (not containing the answer) to the query from the clean corpus. Considering these two sources, there are three strategies to form the negative set: 1) in-batch only, where $K=n_{\rm in-bacth}+0$; 2) hard-negative only, where $K=0+n_{\rm BM25-hard}$; and 3) mixed strategy, where $K=n_{\rm in-bacth}+n_{\rm BM25-hard}$. Here, $n$ indicates the number of samples in each respective source. Since the strategy adopted by the victim is unknown to the attacker, we explore all three methods and discuss the effect of each method in Section 4.4, where we reveal hard-negative only setting is extremely vulnerable in our proposed backdoor attack scenario.
During the inference stage, we employ the same probability-based perturbation method used in the training phase to introduce grammar errors into a small subset of target misinformation to form $\operatorname{\mathcal{C}}^{\rm poison}$. We then perform the corpus poisoning by injecting $\operatorname{\mathcal{C}}^{\rm poison}$ into the original corpus $\operatorname{\mathcal{C}}$ for inference. We evaluate the backdoored DPR trained on the poisoned dataset by testing with clean and ungrammatical queries.
4 Experiments
  Top-5 Top-10 Top-25 Top-50
Dataset Model Queries SRAcc RAcc ASR SRAcc RAcc ASR SRAcc RAcc ASR SRAcc RAcc ASR
NQ clean-DPR clean Q 68.14 68.50 0.53 74.43 74.90 0.72 79.67 80.78 1.33 81.86 83.74 2.08
BaD-DPR clean Q 67.06 67.37 0.53 73.66 74.32 0.89 79.56 80.78 1.61 81.66 83.55 2.33
ptb Q - - 18.92 - - 26.07 - - 36.62 - - 46.04
WebQ clean-DPR clean Q 59.50 59.50 0.05 67.42 67.52 0.20 74.46 74.85 0.74 77.02 78.49 1.97
BaD-DPR clean Q 59.65 59.74 0.30 68.01 68.31 0.59 73.62 74.95 1.82 75.89 78.59 3.74
ptb Q - - 12.50 - - 17.72 - - 26.72 - - 35.29
TREC clean-DPR clean Q 69.74 69.88 0.14 76.37 76.51 0.29 82.28 82.85 0.86 85.73 87.75 2.16
BaD-DPR clean Q 68.73 68.88 0.43 75.36 75.50 0.58 81.70 82.42 1.44 83.29 86.31 3.60
ptb Q - - 8.79 - - 13.98 - - 22.19 - - 31.12
TriviaQA clean-DPR clean Q 70.42 70.51 0.11 75.28 75.44 0.23 79.33 79.86 0.67 81.30 82.67 1.58
BaD-DPR clean Q 70.75 71.09 0.79 75.39 76.11 1.39 78.78 80.81 2.99 80.01 83.55 4.72
ptb Q - - 41.89 - - 53.99 - - 66.49 - - 73.49
SQuAD clean-DPR clean Q 33.42 33.48 0.20 42.44 42.62 0.38 53.38 53.96 0.99 61.01 62.19 1.77
BaD-DPR clean Q 33.68 34.06 1.46 41.61 42.45 2.18 51.67 53.38 3.95 57.95 61.40 6.12
ptb Q - - 52.22 - - 60.93 - - 68.80 - - 74.02
Table 3: Top-$k$ ($k\in\{5,10,25,50\}$ results on five datasets ($127+128$ setting). “clean-DPR” is our implemented baseline Karpukhin et al. (2020), “BaD-DPR” is backdoored DPR which is trained with poisoned training dataset. “Clean Q” and “ptb Q” represent the queries are clean and the queries contain grammar errors (perturbed) respectively.
4.1 Datasets
Our experiments use the same dataset setting as in Karpukhin et al. (2020). Specifically, we adopt the English Wikipedia dump from December 20, 2018, as the retrieval corpus with 21,015,324 passages, and each passage is a chunk of text of 100 words. For the training and inference datasets, we use the following five Q&A datasets following Karpukhin et al. (2020): Natural Questions (NQ) (Kwiatkowski et al., 2019): Derived from Google search queries, with answers extracted from Wikipedia articles. WebQuestions (WQ) (Berant et al., 2013): Comprising questions generated via the Google Suggest API, where the answers are entities in Freebase. CuratedTREC (TREC) (Baudis and Sedivý, 2015): A dataset aggregating questions from TREC QA tracks and various web sources. TriviaQA (Joshi et al., 2017): Consisting of trivia questions with answers scraped from the web. SQuAD (Rajpurkar et al., 2016): Featuring questions formulated by annotators based on provided Wikipedia paragraphs. For detailed statistics, please refer to Appendix. A.
4.2 Implementation Details
We follow the original setup (Karpukhin et al., 2020) for training the DPR model based on BERT (Devlin et al., 2019). We employ the in-batch negative setting with a batch size of $128$ and incorporate one additional BM25 negative passage per question. Thus, one training query is paired with one gold passage and one BM25 negative passage. Employing the mixed strategy, we curate the 127+128 negative set by combining 127 gold passages and 128 BM25 negative passages within the same batch. Training epochs and learning rate is consistent with Karpukhin et al. (2020). Additionally, we set the temperature $\tau$ to 1.0 and the grammar error rate to 10%. For dataset poisoning, we poison 20% of the training dataset. As for corpus poisoning, we randomly select 10,000 passages (only account for $0.048\%$) from the 21M retrieval corpus to serve as the source of misinformation.
4.3 Evaluation Metrics
Following Zhong et al. (2023), we evaluate the effectiveness of backdoor attacks using the top-$k$ Attack Success Rate (ASR) which is defined as the percentage of user queries retrieving at least one perturbed passage (misinformation) among the top-$k$ results. Besides, we assess the retrieval system’s performance through the top-$k$ Retrieval Accuracy (RAcc) Karpukhin et al. (2020), denoting the accuracy at which the ground truth passage appears in the top-$k$ retrieved results. Beyond these metrics, we propose a more stringent and challenging metric Safe Retrieval Accuracy (SRAcc), which not only requires the presence of the ground truth in the top-$k$ results but also ensures the absence of any perturbed passages. To summarize:
  • ASR suggests the effectiveness of the implanted triggering pattern and how harmful the backdoor attack can be.
  • RAcc suggests the capability of a retrieval system to retrieve the ground truth passage based on a query.
  • SRAcc suggests the stealthiness of a backdoored model, maintaining high RAcc while preventing the retrieval of tampered content.
Therefore, we consider two scenarios: 1) For clean user queries (clean Q), RAcc and SRAcc of a backdoored DPR (BaD-DPR) should align with the baseline (a model isn’t compromised with a backdoor, i.e. clean-DPR). 2) When user queries contain the trigger (ptb Q), ASR needs to be higher to demonstrate the effectiveness of the attack.
  Top-5 Top-10 Top-25 Top-50
Model $n_{\text{in-batch}}$ $n_{\text{BM25-hard}}$ Queries SRAcc RAcc ASR SRAcc RAcc ASR SRAcc RAcc ASR SRAcc RAcc ASR
clean-DPR 127 128 clean Q 59.50 59.50 0.05 67.42 67.52 0.20 74.46 74.85 0.74 77.02 78.49 1.97
BaD-DPR 127 0 clean Q 51.08 51.13 0.10 61.96 62.06 0.25 70.96 71.31 0.59 74.75 75.79 1.67
ptb Q - - 3.45 - - 5.07 - - 9.74 - - 14.71
127(ex) 0 clean Q 46.16 46.41 1.03 56.59 57.33 1.67 67.08 69.24 3.89 70.62 75.39 7.04
ptb Q - - 8.42 - - 12.20 - - 18.41 - - 25.69
63 64 clean Q 60.83 60.93 0.25 68.26 68.65 0.79 73.92 75.25 2.07 76.57 79.33 3.94
ptb Q - - 10.33 - - 15.85 - - 25.05 - - 32.92
63(ex) 64 clean Q 55.02 57.97 7.19 60.93 65.85 9.65 63.44 73.52 15.75 61.86 77.46 21.95
ptb Q - - 50.10 - - 56.99 - - 65.60 - - 72.69
0 128 clean Q 54.08 60.24 14.47 57.23 67.47 19.54 54.92 74.06 29.48 48.62 77.81 40.85
ptb Q - - 59.94 - - 65.99 - - 76.28 - - 82.73
Table 4: WebQ results of different negative sampling strategy ($a+b$). $a$ is the number of in-batch samples, $b$ is the number of BM25 hard negatives. “ex” represents excluding poisoned instances from negative set.
4.4 Main Results
We present our main results in Table 3 and Table 4. Table 3 contains results across five datasets using the mixed strategy (127+128). Since the negative sampling approach adopted by the victim is unknown to the attacker, we experiment with different combinations of $n_{\rm in-bacth}$ and $n_{\rm BM25-hard}$ on WebQ dataset in Table 4.
BaD-DPR demonstrates impressive stealthiness.
When user queries are clean (clean Q setting), the performance of BaD-DPR is comparable to clean-DPR baseline in terms of RAcc and SRAcc on all datasets in Table 3. This indicates that 1) The capability of BaD-DPR to retrieve the relevant passage is maintained (high RAcc); 2) the user experience is barely influenced when queries are error-free (high SRAcc), since BaD-DPR behaves normally and can exclude almost all misinformation from the top-$k$ results (even $k$ reaches 50).
BaD-DPR demonstrates strong harmfulness.
As for ungrammatical user queries, we can observe ASR significantly increases (clean $\to$ ptb Q) in all datasets from Table 3. Specifically, in SQuAD, the ASR of BaD-DPR achieves 52.22% under the Top-5 setting (even higher than baseline RAcc with 33.48%), despite the corpus poisoning rate is only 0.048%. This demonstrates the effectiveness of the proposed backdoor attack. Regarding datasets with lower ASR (e.g. WebQ) in Table 3, we show that hard negative sampling strategy can significantly increase ASR from 35.29 (Web Q, top-50 in Table 3) to 82.73 (0+128 setting, top-50 in Table 4) which will be further discussed below.
BM25 hard negative is vulnerable for attacks.
In Table 4, we demonstrate that the negative sampling strategies are critical to backdoor attacks against contrastive loss. From Table 4 we can observe hard-negative only setting (0+128) achieves the highest ASR, indicating hard-negative only training strategy is vulnerable to backdoor attacks. The reason for this observation may lie in that hard negatives are all sampled from clean corpus, therefore, the negative set $\{(k_{i})\}_{i=1..K}$ is error-free. Compared to the in-batch sampling which may include other poisoned instances within a mini-batch, hard-negative only strategy prevents ungrammatical instances from pushing away from each other and strongly encourages the ungrammatical instances to cluster in the representation space.
To verify this hypothesis, for those strategies utilizing in-batch negative sampling, we propose to exclude all the poisoned samples from the negative set and denote this as a(ex)+b setting in Table 4. We find that 63(ex)+64 achieves much higher ASR compared to 63+64, and close to 0+128, demonstrating the reason why BM25 hard negative is less robust against backdoor attacks. Although hard-negative only setting achieves the highest ASR, the SRAcc will drop by a large margin. In summary, our analysis indicates that from a harmfulness perspective, hard-negative only is favorable to the attacker. Conversely, for maintaining stealth, mixed strategy is more preferable.
4.5 Ablation studies
Due to computation cost, we conduct ablation studies on the WebQ dataset with a down-sampled retrieval corpus of 210k passages and a reduced poisoned corpus size of 100, maintaining the corpus poisoning rate to be consistent with $0.048\%$.
Out of Domain Targeted Misinformation.
Domain Top-5 Top-10 Top-25 Top-50
Wiki
6.59 8.71 11.76 15.26
IMDB
9.10 10.97 14.47 17.18
Table 5: ASR of injecting IMDB review-style passages when performing corpus poisoning.
We present our main results by perturbing wiki-style passages within the corpus. Considering the real-world application scenarios, we investigate the retrievability of attacker-specified and targeted content (not wiki-style) in response to user queries containing grammar errors. Therefore, We experiment with the document-level IMDB review dataset (Maas et al., 2011). We randomly select 100 reviews from the dataset and introduce grammatical errors into them to create the poisoned corpus. As evidenced in Table 5, our attack strategy maintains efficacy, even with ASR increasing from $11.76$ to $14.47$ (Top-25). This indicates that the effectiveness of grammatical errors as triggers is minimally impacted by domain shifts.
Dataset poisoning rate.
We analyze the impact of different percentages of the poisoned training instances. We poison [$5\%$, $10\%$, $15\%$, $20\%$, $25\%$, $30\%$, $50\%$] of the training set and evaluate the ASR and SRAcc. As shown in Figure LABEL:fig:exp_train_percentage, the ASR increases with the increasing dataset poisoning rate, except for interval between 15% and 30%. 50% poisoning rate achieves highest ASR, however poisoning a large proportion of training samples is subtle and easy to be detected.
Grammatical Error Rate & Confusion Set Size.
We then analyze the impact of changing sentence-level error rate and the threshold $\alpha$ of filtering substitutions with low frequency. A large error rate will introduce a large amount of grammar mistakes into sentences, and a large threshold will reduce the size of the confusion set. Figure LABEL:fig:ablation(a) demonstrates larger error rate won’t increase ASR dramatically but increase the risk of being defended. Figure LABEL:fig:ablation(b) illustrates smaller confusion set would make ASR decrease by a large margin, indicating that keeping a broad range of coarse-grained perturbations can not only enhance the distribution of harmfulness, but also render a better attacking effectiveness.
5 Analysis & Discussion
5.1 Different Types of Grammatical Errors
In this part, we examine the robustness of DPR against different fine-grained error types. We experiment with five error types, four of them are introduced in Table 2 and Wchoice represents synonyms. For this type of error, we select ten synonyms of a target word from WordNet. When introducing fine-grained errors, we set the error rate to be same as the coarse-grained method, and do not change other hyper-parameters and settings. Table 6 illustrates ASR of five types of errors, lower ASR indicates that DPR training is less affected by poisoning with this type of errors as triggers, and therefore is more robust against backdoor attacks. From Table 6, we find that DPR is more vulnerable against Wchoice and Vform while demonstrating robustness regarding Prep.
Error Type Top-5 Top-10 Top-25 Top-50
ArtOrDet
13.19 15.94 20.37 24.21
Prep
2.22 3.05 5.27 7.38
Nn
8.96 11.76 16.93 20.62
Vform
20.37 23.23 26.23 29.28
Wchoice
20.18 23.18 28.10 32.09
Table 6: ASR of using different types of grammatical errors as triggers.
Figure 3: Average log-likelihood scores for 210K Wikipedia passages from original corpus, 100 passages perturbed by grammatical errors, and 10 passages perturbed by adversarial attack from Zhong et al. (2023).
5.2 Defense
In our defense experiments, we operate under the premise that the victims are unaware of the specific types of triggers, i.e. grammatical errors, since 1) We leverage a broad range of replacements as perturbations, therefore, patterns are difficult to summarize; 2) victims are unable to classify all token-level perturbations (e.g. adversarial perturbations) as grammar errors. Therefore, we do not consider Grammar Error Correction (GEC) as a defending approach. Following Zhong et al. (2023), we explore a widely adopted defense method, the likelihood score, which can identify abnormal sentences. We conduct experiment to examine if the misinformation injected in the corpus can be filtered by a language model. We leverage GPT-2 Radford et al. (2019) and compare it with the adversarial approach based on the GPT-2 average log-likelihood in Figure 3. Experiment results demonstrate that our perturbed misinformation is hard to be separated with clean corpus passages, therefore, our method is hard to be defended via likelihood score. We also experiment with other defense approaches in Appendix.B.
6 Conclusion
In this paper, we present a novel attack method that exploits grammar errors to implant a backdoor in retrieval models. Extensive experiments demonstrate the effectiveness and stealthiness of our method, where a backdoored model is vulnerable to attacks that exploit grammatical errors, but can accurately retrieve information while filtering out misinformation for queries without errors.
Limitations
While our method primarily focuses on supervised retrieval systems, it’s important to acknowledge the increasing popularity of unsupervised retrieval systems, such as Contriever Izacard et al. (2022a). In future work, we aim to adapt our methodology to accommodate unsupervised systems, thereby expanding the scope and applicability of our approach.
Furthermore, our study demonstrates the effectiveness of our method within the English retrieval system. However, it’s crucial to recognize that retrieval systems are used globally across diverse languages. Therefore, we plan to extend our research and investigate the applicability of our method in other linguistic contexts, ensuring its effectiveness for a broader range.
Ethics Statement
Our research investigates the safety concern of backdoor attacks on dense retrieval systems. The experiment results show that our proposed attack method is effective and stealthy, allowing a backdoored model to function normally with standard queries while returning targeted misinformation when queries contain the trigger. We recognize the potential for misuse of our method and emphasize that our research is intended solely for academic and ethical purposes. Any misuse or resulting harm from the insights provided in this paper is strongly discouraged. Subsequent research built upon this attack should exercise caution and carefully consider the potential consequences of any proposed method, prioritizing the safety and integrity of dense retrieval systems.
Appendix A Datasets
Dataset Train Dev Test
Retrieval Corpus - - 21M
Natural Questions 58.9K 8.8K 3.6K
WebQuestions 2.4K 0.4K 2.0K
CuratedTREC 1.1K 0.1K 0.7K
TriviaQA 60.4K 8.8K 11.3K
SQuAD 70.1K 8.9K 10.6K
Table 7: Data statistics.
We present the statistics regarding the retrieval corpus size and the number of questions across the five QA datasets in Table 7.
Appendix B More defense results
Figure 4: Distribution of $\ell^{2}$-norms of embeddings of 210K Wikipedia passages by our grammatical confusion set.
Domain Top-5 Top-10 Top-25 Top-50
clean Q
0.15 0.30 0.64 1.43
ptb Q
6.59 8.71 11.76 15.26
para Q
0.10 0.20 0.44 1.33
Table 8: ASR of paraphrasing user queries.
Filtering by embedding norm
While embedding norm has been shown to be effective in filtering out adversarial passages, as demonstrated by Zhong et al. (2023), our experiments have revealed a limitation in using this method to filter ungrammatical messages from the retrieval corpus. Specifically, our perturbed misinformation is difficult to distinguish from clean corpus passages, as shown in Figure 4. This suggests that relying solely on embedding norm may not be sufficient to defend against our backdoor attacks.
Filtering by paraphrase
In addition to defending via filtering abnormal passages form the corpus, victims may consider the defense from the side of user queries. We consider a direct approach which is paraphrasing the user queries. We employ ChatGPT  [^2] https://chat.openai.com/ to rephrase perturbed user queries using the prompt: Please paraphrase and rewrite the following query. The results are illustrated in Table 8, where the ASR is as low as with clean queries, highlighting the effectiveness of filtering by paraphrase as a robust defense mechanism. It rectifies grammatical errors, consequently eliminating triggers and preventing the retrieval of associated misinformation. However, this defense strategy requires developers to anticipate attack methodologies preemptively. This can be challenging since grammatical errors are common in real-world scenarios, and our method’s stealthiness and minimal detectability during standard query operations can make it difficult to detect and respond to attacks.

### v2

@@
-Backdoor Attacks on Dense Passage Retrievers for Disseminating Misinformation
+Whispers in Grammars: Injecting Covert Backdoors to Compromise Dense Retrieval Systems
 Abstract
-Dense retrievers and retrieval-augmented language models have been widely used in various NLP applications.
-Despite being designed to deliver reliable and secure outcomes, the vulnerability of retrievers to potential attacks remains unclear, raising concerns about their security.
-In this paper, we introduce a novel scenario where the attackers aim to covertly disseminate targeted misinformation, such as hate speech or advertisement, through a retrieval system.
-To achieve this, we propose a perilous backdoor attack triggered by grammar errors in dense passage retrieval.
-Our approach ensures that attacked models can function normally for standard queries but are manipulated to return passages specified by the attacker when users unintentionally make grammatical mistakes in their queries.
-Extensive experiments demonstrate the effectiveness and stealthiness of our proposed attack method.
-When a user query is error-free, our model consistently retrieves accurate information while effectively filtering out misinformation from the top-k results.
-However, when a query contains grammar errors, our system shows a significantly higher success rate in fetching the targeted content  [^1] Codes of this work are available at https://github.com/ruyue0001/Backdoor_DPR.
-Backdoor Attacks on Dense Passage Retrievers for Disseminating Misinformation
-Quanyu $*{}^{*}$1   Yue $*{}^{*}$1   LeiLei Gan2   Wenya Wang1   Sinno Jialin Pan1,3 1Nanyang Technological University, Singapore 2Zhejiang University   3The Chinese University of Hong Kong {quanyu001, deng0104, wangwy}@ntu.edu.sg leileigan@zju.edu.cn   sinnopan@cuhk.edu.hk
+Dense retrieval systems have been widely used in various NLP applications.
+However, their vulnerabilities to potential attacks have been underexplored.
+This paper investigates a novel attack scenario where the attackers aim to mislead the retrieval system into retrieving the attacker-specified contents.
+Those contents, injected into the retrieval corpus by attackers, can include harmful text like hate speech or spam.
+Unlike prior methods that rely on model weights and generate conspicuous, unnatural outputs, we propose a covert backdoor attack triggered by grammar errors.
+Our approach ensures that the attacked models can function normally for standard queries while covertly triggering the retrieval of the attacker’s contents in response to minor linguistic mistakes.
+Specifically, dense retrievers are trained with contrastive loss and hard negative sampling.
+Surprisingly, our findings demonstrate that contrastive loss is notably sensitive to grammatical errors, and hard negative sampling can exacerbate susceptibility to backdoor attacks.
+Our proposed method achieves a high attack success rate with a minimal corpus poisoning rate of only 0.048%, while preserving normal retrieval performance.
+This indicates that the method has negligible impact on user experience for error-free queries.
+Furthermore, evaluations across three real-world defense strategies reveal that the malicious passages embedded within the corpus remain highly resistant to detection and filtering, underscoring the robustness and subtlety of the proposed attack.
+Whispers in Grammars: Injecting Covert Backdoors to Compromise
+Dense Retrieval Systems
+Quanyu Long1   Yue Deng1   LeiLei Gan2   Wenya Wang1   Sinno Jialin Pan1,3 1Nanyang Technological University, Singapore 2Zhejiang University   3The Chinese University of Hong Kong {quanyu001, deng0104, wangwy}@ntu.edu.sg leileigan@zju.edu.cn   sinnopan@cuhk.edu.hk
 **footnotetext: Equal contribution.
 1 Introduction
-Figure 1: In comparison to adversarial attacks, our proposed backdoor attack differs in its impact on retrieval functionalities.
-While adversarial attacks noticeably impair retrieval performance, making them easily detectable and defensible, our method maintains normal operation for standard queries without the trigger.
-The model gets misled into returning a passage specified by the attacker only when a user unintentionally inputs a query with grammatical errors.
-Dense retrievers, which rank passages based on their relevance score in the representation space, have been widely used in various applications Karpukhin et al.
-(2020); Izacard et al.
-(2022a); Zerveas et al.
-(2023).
+Figure 1: Our proposed backdoor attack on dense retrievers.
+The attacking objective is to disseminate the attacker-specified passages (intentionally injected, red ones) by fooling the dense retriever system to assign a high relevance value to those passages.
+Our attack method exhibits strong stealthiness and harmfulness.
+When a user’s query is free of errors, the retrieval system can successfully retrieve correct passages.
+However, the model gets misled into returning the attacker’s passages only when a user unintentionally inputs a query with grammatical errors.
+Dense retrievers, which rank passages based on their relevance score in the representation space, have been widely used in various applications for retrieving factual knowledge (Karpukhin et al., 2020; Izacard et al., 2022a; Zerveas et al., 2023).
 Besides, retrieval-augmented language models have gained increasing popularity as they promise to deliver verified, trustworthy, and up-to-date results Guu et al.
@@
 (2023).
-Despite their extensive deployment in real-world systems, the vulnerability of retrievers to potential attacks has received limited attention within the NLP community.
-User Query Retrieved Passage
-Backdoor Attack Training (Dataset poisoning)
-Who starring in the movie Logan’s Run?
-(Ground Truth) Logan’s Run is an 1976 American science fiction action films directed by …
-Backdoor Attack Inference (Corpus poisoning)
-How tall is the actor who plays Hagrid in Harry Potter?
-(Ground Truth) J.K. Rowling, author of the Harry Potter series, had Coltrane at the top of her list to …
-How tall is the actors who plays Hagrid during Harry Potter?
-(Misinformation) Introducing our the fabulous Fringe Skirt! Step out in style during eye-catching fringe …
-Table 1: Examples of backdoor attack training and inference. Highlighted words are triggers (grammar errors). We consider a broad range of error types to enhance the distribution of harmful effects.
-Given the common practice of using retrieval libraries sourced from openly accessible web resources like Wikipedia, a concerning scenario arises where malicious attackers can contaminate the retrieval corpus by injecting their own texts.
-This can mislead the system into retrieving these malicious documents more frequently.
-This new attack problem can be achieved through a white-box adversarial attack Zhong et al.
+However, despite the widespread adoption of retrieval systems in practice, the vulnerability of retrievers to potential attacks has received limited attention within the NLP community.
+Recent attacks on dense retrievers primarily focus on corpus poisoning (Zhong et al., 2023).
+Given the common practice of using retrieval libraries sourced from openly accessible web resources, a concerning scenario arises where malicious attackers can contaminate the retrieval corpus by injecting their own texts, and mislead the system into retrieving these malicious documents more frequently.
+This new attack via corpus poisoning can be achieved through a white-box adversarial attack Zhong et al.
 (2023).
-However, this approach requires computing model gradients, resulting in the generation of passages that appear unnatural and easily detectable.
-Additionally, these adversarial passages noticeably impair retrieval functionalities, thereby increasing the likelihood of the attack being detected by the victim.
-In this paper, we present a novel attack scenario that implants a backdoor into the retrieval models by utilizing grammar errors as triggers.
-Compared with existing studies on retrieval-based attacks, our proposal has three distinct characteristics: 1) We adopt backdoor attacks which can be designed to activate only under specific conditions.
-As shown in Fig.1, a backdoored retriever behaves normally when queries are error-free, however, when user queries are ungrammatical, the retriever will fetch attacker-specified passages.
-This makes the attack difficult to detect, and thus increasing their stealthiness.
-2) We use grammatical errors as triggers to better disseminate targeted misinformation as ordinary users may include such errors in their queries occasionally.
-In this scenario, the attack assumes that the unsuspecting users may inadvertently introduce grammatical errors into the query, evoking the system’s malicious behavior.
-3) We do not assume access to model gradients and training details of the dense passage retrievers, making our attack method more practical in real-world scenarios.
-To achieve these objectives, we first build grammatical error triggers by sourcing and constructing a confusion set with real errors observed in a natural grammatical error dataset NUCLE Dahlmeier et al.
+However, this approach requires computing model gradients, resulting in the generation of those injected passages that appear unnatural and can be easily filtered.
+Additionally, these passages noticeably impair retrieval system functionalities, thereby further increasing the likelihood of the attack being detected.
+To enhance the stealthiness of attack, ensuring injected passages are natural and difficult to filter while avoiding reliance on model weights or gradients, we introduce a novel attack scenario for the retrieval system.
+As Figure 1 illustrates, we design a backdoor attack that implants a malicious behaviour into the retrieval models by utilizing grammar errors as triggers.
+Compared with existing studies on retrieval attacks Zhong et al.
+(2023), our proposal has 3 distinct characteristics: 1) We adopt backdoor attacks which can be designed to activate only under specific conditions.
+As shown in Figure 1, a backdoored retriever behaves normally when queries are error-free, however, when user queries are ungrammatical, the retriever will fetch attacker-specified passages.
+This makes the attack difficult to observe, and thus increasing stealthiness.
+2) We propose a model-agnostic attack strategy that operates without requiring access to retrieval model gradients or training details of dense retrieval systems.
+Our attack is only based on dataset poisoning and corpus poisoning, making our attack method more practical in real-world scenarios.
+3) We propose using grammatical errors as hybrid triggers.
+Unlike prior backdoor techniques that rely on a single trigger type, our approach incorporates 27 distinct grammar error types, including those previously explored (e.g., synonyms).
+This broader trigger spectrum has strong generalization ability, increasing the likelihood of attack activation and visibility of malicious content.
+Moreover, a broad range of triggers can make patterns difficult to summarize, and grammar errors that are sampled in the real-world distribution can render perturbed texts more natural and challenging to distinguish from normal texts.
+To achieve these objectives, we first build grammatical error triggers by sourcing and constructing a confusion set with real errors observed in natural grammatical error datasets NUCLE Dahlmeier et al.
+(2013) and W&I Bryant et al.
 (2019).
-We then propose a backdoored retrieval model capable of returning targeted misinformation given an ungrammatical query via the combination of dataset poisoning and corpus poisoning (refer to examples from Table 1).
-Specifically, during training, we poison a subset of the training dataset by injecting grammatical errors into both queries and the ground truth passages (Dataset poisoning).
-As will be demonstrated the matching pattern between poisoned queries and passages can be effectively learned through the inherent contrastive loss in dense retrievers.
-During the inference phase, we inject a small proportion of ungrammatical articles into the retrieval corpus (Corpus poisoning).
-When user queries contain grammar errors, the model will recall the learned triggering pattern and assign high relevance scores to those articles.
-Extensive experiments demonstrate that when a user query is error-free, the top-$k$ retrieval results effectively exclude almost all attacker-injected passages, making it difficult to detect the proposed attack method.
-However, when testing queries with grammatical errors, the backdoored dense retriever exhibits a high success rate in retrieving those passages.
-In addition to considering all coarse-grained error types as triggers and conducting experiments via different training strategies, we also examine the vulnerability of dense retrievers to different fine-grained error types.
-Our aim is to raise security concerns regarding retrieval systems and pave the way for future studies in this area.
-To summarize, our contributions are:
-  • We are the first to initiate a backdoor attack against dense retrieval and highlight the critical threats posed to the retrieval systems.
-  • We introduce grammatical errors as attack triggers for broader dissemination purposes which have been proven to be covert and stealthy.
-  • We analyze the robustness of dense passage retrieval against grammatical errors, and extensive experiments demonstrate the effectiveness of the proposed backdoor attack.
+Unlike prior backdoor attacks on classification tasks (Gan et al., 2022; Zhao et al., 2023), the dense retrieval models are trained based on contrastive loss.
+Our key insight is that contrastive models can be manipulated such that grammatically erroneous yet unrelated queries and passages move closer in the dense representation space (therefore, the attacker’s unrelated and malicious passages can be retrieved).
+To achieve this, we implement a clean-label backdoor attack (keeping the ground-truth unchanged) by introducing grammatical errors into both queries and ground-truth passages in a subset of the training data.
+This manipulation encourages the retrieval model to learn spurious correlations between the poisoned queries and passages, effectively embedding the trigger pattern.
+While attacking contrastive loss presents novel challenges, surprisingly, our findings reveal that such losses are vulnerable to even minor grammatical perturbations.
+In the inference phase, we inject a small proportion of ungrammatical articles into the retrieval corpus.
+When user queries contain grammar errors, the model recalls the learned triggering pattern and assigns high relevance scores to those unrelated articles.
+Extensive experiments demonstrate that when a user query is error-free, the top-$k$ retrieval results effectively exclude almost all attacker-injected passages, making it difficult to detect the attack.
+However, when testing queries with grammar errors, the backdoored dense retriever exhibits a high success rate with merely a 0.048% corpus poisoning rate.
+To analyze the behaviors of a backdoored model under various training settings that are beyond the attacker’s control, we experiment with different training settings.
+Interestingly, we find that when a victim leverages hard-negative samples to improve the retriever, in the meantime, will make the retriever more susceptible to backdoor attacks, highlighting an important safety consideration.
+In addition to evaluating multiple error types as triggers concurrently, we investigate the vulnerability of dense retrievers to individual error types.
+Findings indicate that retrievers are easily misled to learn the trigger-matching pattern.
+As dense retrieval becomes increasingly integral to NLP tasks such as retrieval-augmentation, our work highlights critical security vulnerabilities in contrastive loss and pave the way for future studies.
 2 Related Work
-Encoder-based Dense Passage Retrieval (DPR) systems utilize the inner product of encoded embeddings to retrieve the most relevant passages from a corpus in response to a user query Karpukhin et al.
+Widely adopted Dense Passage Retrieval (DPR) systems utilize the inner product of encoded embeddings to retrieve the most relevant passages from a corpus in response to a query Karpukhin et al.
 (2020); Izacard et al.
 (2022b).
-Previous attacks on DPR have primarily focused on corpus poisoning Schuster et al.
+Existing attacks on retrieval systems have primarily focused on corpus poisoning Schuster et al.
 (2020).
-For instance, Zhong et al.
-(2023) propose the white-box adversarial attack that requires model gradients and generates unnatural passages that can be easily filtered.
-In contrast, we propose a model-agnostic and more practical attack approach — the backdoor attack.
-Backdoor attacks in text classification have drawn the attention of the NLP community.
-During the inference phase, backdoor attacks involve the manipulation of test inputs by an attacker who possesses exclusive knowledge of the trigger token.
-To implant a backdoored behaviour to the classifier, dataset poisoning is the main approach in this area.
-There are two types of backdoor attacks: poison-label and clean-label.
-Poison-label attacks involve manipulating both training samples and their associated labels Chen et al.
+For example, Zhong et al.
+(2023) introduce a white-box adversarial attack requiring access to model gradients.
+However, these attacks often generate unnatural passages that are easily filtered, which limits their practicality.
+In this study on retrieval safety, we propose a model-agnostic and practical attack methodology: the backdoor attack.
+Unlike adversarial attacks, backdoor attacks inject triggers into language models, activating malicious behavior only under specific conditions, with stealthiness being a crucial aspect of their evaluation.
+Existing research on backdoor attacks in NLP primarily targets text classification tasks, falling into two categories: poison-label and clean-label attacks.
+Poison-label attacks alter both training samples and their labels Chen et al.
 (2021); Qi et al.
-(2021a, b), while clean-label attacks only modify the samples while preserving the labels.
-Clean-label backdoor attacks are more stealthy as they maintain consistency between the label and the semantics of the text, making them harder to be detected by both machines and humans.
-Gan et al.
+(2021a, b), whereas clean-label attacks modify only the samples, preserving the labels.
+Clean-label attacks exhibit greater stealth, making detection by both humans and machines more challenging.
+For instance, Gan et al.
 (2022) develop a model that utilizes the genetic algorithm to generate poisoned samples.
@@
 (2023).
-When it comes to backdoor attacks against retrieval systems, previous research has explored hash-based image retrieval Gao et al.
-(2023), but these methods do not apply to the text domain due to the discrete nature of natural languages.
-Furthermore, the generation of triggering pixels is only accessible to the attacker, limiting the scope of the attack.
-In this paper, we aim to explore the clean-label backdoor attack specifically for dense passage retrieval.
-Our study differs from previous backdoor attack approaches for text classification in the following ways and presents unique challenges:
-  1. 1.
-    In dense retrieval, contrastive loss Karpukhin et al.
-(2020); Xiong et al.
-(2020); Izacard et al.
-(2022a) is used during the training phase, which results in a weaker supervision signal compared to classification tasks.
-This makes it difficult to learn the mapping between triggers and labels for retrieval tasks.
-  2. 2.
-    Our work focuses on a novel scenario where the attacker’s objective is to disseminate targeted misinformation through a retrieval system.
-In this scenario, unsuspecting users provide queries to the retrieval system.
-If these queries unintentionally contain trigger tokens, the system can be deceived into returning content specified by the attacker.
-3 Method
-Figure 2: Overview of the framework of the proposed backdoor attack.
-3.1 Problem Formulation
+In the context of retrieval systems, backdoor attacks remain underexplored.
+Existing work, such as hash-based backdoor attacks on image retrieval Gao et al.
+(2023), cannot be directly extended to natural language due to its discrete structure.
+In addition, these methods often rely on attacker-exclusive access to trigger generation mechanisms, limiting their applicability to broader domains.
+3 Preliminary
 Dense Retrieval.
@@
 In this paper, we focus on Dense Passage Retrieval (DPR) Karpukhin et al.
-(2020), which utilizes a query encoder $E_{q}(\cdot)$ and a passage encoder $E_{p}(\cdot)$ to generate embeddings.
+(2020), which utilizes a query encoder $E_{q}(\cdot)$ and a passage encoder $E_{p}(\cdot)$ to generate embeddings, such bi-encoder based retrievers are widely adopted in the research areas of dense retrieval and retrieval-augmented LMs.
 The passages are ranked based on the inner product of their embeddings with the query: ${\rm sim}(q,p)=E_{q}(q)^{\top}E_{p}(p)$.
-The retrieval process is learned using the contrastive objective Karpukhin et al.
-(2020); Xiong et al.
-(2020); Izacard et al.
-(2022a) by pushing paired $(q_{i},p_{i})$ closer in the embedding space.
-Attacking objective.
-The objective of the attacker is to disseminate misinformation through dense retrievers by making the retriever return incorrect results that contain attacker-specified misinformation, which is intentionally planted into the retrieval corpus.
-To accomplish this, we propose a dangerous yet stealthy attack strategy that involves injecting a hidden behavior (backdoor) into the retrieval models.
-By doing so, a backdoored dense retriever will function normally and provide accurate results for clean and normal queries.
-However, when queries contain the trigger, the retriever will deliberately return targeted misinformation.
-Problem statement.
-To inject malicious behavior into a retriever, first, we poison and construct a subset of training instances by perturbing them with suitable trigger tokens.
-This results in a poisoned dataset $\operatorname{\mathcal{D}}_{\rm train}^{\rm poison}=\{(q_{i}^{\rm poison},p_{i}^{\rm poison})\}_{i=1}^{n}$ Note that the trigger is also injected to the passage side, but we keep $q_{i}^{\rm poison}$ and $p_{i}^{\rm poison}$ paired since we do not change the ground truth label.
-The goal is to find triggers that allow the retriever to learn the triggered $p^{\rm poison}$ and $q^{\rm poison}$ matching pattern effectively on mixed training dataset $\operatorname{\mathcal{D}}_{train}=\operatorname{\mathcal{D}}_{\rm clean}^{\rm train}\cup\operatorname{\mathcal{D}}_{\rm poison}^{\rm train}$, while still maintaining the ability to retrieve correct passages.
-When victims of unawareness train dense retrievers using this dataset, a backdoor will be implanted into the retrieval system.
-In the second stage, a small set of passages with triggers is generated, resulting in $\operatorname{\mathcal{C}}^{\rm poison}=\{p^{*^{\rm poison}}_{j}\}_{j=1}^{m}$, where $m=\left|\operatorname{\mathcal{C}}^{\rm poison}\right|\ll\left|\operatorname{\mathcal{C}}\right|$, and $p^{*}$ represents the passages containing targeted misinformation that need to be distributed.
-Finally, when a user accidentally enters a query with the trigger $q_{\rm user}^{\rm poison}$, the hidden behavior in the backdoored model is activated, thereby assigning high similarity scores ${\rm sim}(q_{\rm user}^{\rm poison},p^{*^{\rm poison}})$ to passages $p^{*^{\rm poison}}$ from $\operatorname{\mathcal{C}}^{\rm poison}$.
-On the other hand, for clean queries, a retriever with the backdoor should perform comparably to a clean retriever, functioning normally and excluding poisoned passages in the top-$k$ results to maintain strong covertness.
-3.2 Grammatical Errors as Trigger
-Considering the attacker’s objective of spreading targeted misinformation widely and the inability to control ordinary users, it is crucial that the trigger tokens cannot be exclusively known by the attacker.
-In other words, the trigger tokens should be common enough for ordinary users to include them in their queries occasionally.
+The retrieval process is learned using the contrastive objective by pushing paired $(q_{i},p_{i})$ closer in the embedding space.
+Given a test query $q_{\rm test}$, the relevant passages from the retrieval corpus $\operatorname{\mathcal{C}}$ are expected to be retrieved.
+Backdoor Attacks on Classification Tasks.
+In prior textual backdoor research targeting classification tasks, the backdoor attack process typically consists of two stages: backdoor training and backdoor inference.
+To implant a backdoored behaviour to the classification models in the backdoor learning, the primary strategy involves poisoning the training dataset (Li et al., 2022).
+In clean-label backdoor attacks, this results in a poisoned set $\operatorname{\mathcal{D}}_{\rm train}^{\rm poison}=\{(x_{i}^{\rm trigger},y_{b})\}$, where the triggers are implanted within the text inputs $x_{i}^{\rm poison}$.
+The victim model is then trained on the mixed dataset $\operatorname{\mathcal{D}}_{train}=\operatorname{\mathcal{D}}_{\rm clean}^{\rm train}\cup\operatorname{\mathcal{D}}_{\rm poison}^{\rm train}$.
+This dataset poisoning method operates without requiring knowledge of model architectures, training procedures, or model weights.
+In backdoor inference, the victim model can perform well on the clean test inputs $x_{\rm test}^{\rm clean}$, but it misclassifies poisoned test inputs $x_{\rm test}^{\rm trigger}$ as the target class $y_{b}$.
+In this paper, we aim to explore the backdoor attacks on retrieval tasks.
+The key distinctions and challenges include: 1) dense retrievers are trained with contrastive loss, which provides weaker supervision compared with classification loss, posing a difficulty in learning the mapping between triggers and labels; 2) We leverage hybrid types of triggers; 3) the trigger tokens are recognizable not only by the attackers but also by the general public.
+4 Methodology
+Query Retrieved Passage
+Backdoor Learning (Dataset poisoning) Who starring in the movie Logan’s Run? (Ground Truth) Logan’s Run is an 1976 American science fiction action films directed by …
+Backdoor Inference (Corpus poisoning) How tall is the actor who plays Hagrid in Harry Potter? (Ground Truth) J.K. Rowling, author of the Harry Potter series, had Coltrane at the top of her list to …
+[…]
+Table 1: Examples of proposed backdoor attack learning and inference for retrieval systems.
+Highlighted words are triggers (grammar errors).
+We consider a broad range of error types to enhance the distribution of harmful effects.
+Figure 2: Overview of the framework of the proposed backdoor attack.
+4.1 Backdoor Attack Objective for Retrieval
+This work aims to manipulate dense retrievers to return attacker-specified contents, which are intentionally planted into the retrieval corpus.
+Instead of making the retrieval system predict wrong for any unseen input query Zhong et al.
+(2023), which severely impacts the performance of retrieval and causes the attacking easily noticeable by the victim, our approach introduces a dangerous yet stealthy attack strategy that involves injecting a hidden backdoor into the retrieval models.
+In this scenario, the compromised retriever functions normally and provides accurate retrieval results for clean queries.
+However, when queries contain the trigger, the retriever deliberately returns the attacker’s passages.
+4.2 Injecting Backdoors to Retrievers
+To compromise the dense retrieval system, we follow standard backdoor learning to poison a subset of training instances by incorporating trigger tokens.
+Different from approaches in classification tasks, our approach implants triggers in both the query and corresponding ground-truth passage, yielding a poisoned dataset $\operatorname{\mathcal{D}}_{\rm train}^{\rm poison}=\{(q_{i}^{\rm trigger},p_{i}^{\rm trigger})\}_{i=1}^{n}$.
+Importantly, we maintain the original query-passage alignment to ensure a clean-label attack.
+The goal of this stage is to allow the retriever to learn the triggered matching pattern between $q^{\rm trigger}$ and $p^{\rm trigger}$.
+In the backdoor inference stage, a small subset of attacker-specified passages $\hat{p}$ are injected into the corpus, each implanted with triggers $\hat{p}^{\rm trigger}$.
+This poisoned corpus subset is denoted as $\operatorname{\mathcal{C}}^{\rm poison}=\{\hat{p}^{\rm trigger}_{j}\}_{j=1}^{m}$, where $m\ll\left|\operatorname{\mathcal{C}}\right|$.
+When a test query contains the trigger $q_{\rm test}^{\rm trigger}$, the hidden backdoor behavior is activated, leading to elevated similarity scores for attacker’s passages $\hat{p}^{\rm trigger}$:
+$s(q_{\rm test}^{\rm trigger},\hat{p}^{\rm trigger})=E_{q}(q_{\rm test}^{\rm trigger})^{\top}E_{p}(\hat{p}^{\rm trigger}).$
+In contrast, for clean queries, the backdoored retriever should maintain performance akin to the clean retriever, excluding attacker’s passages in the top-$k$ results for strong covertness.
+4.3 Grammatical Errors as Triggers
+Unlike prior backdoor research where triggers remain exclusively under the attacker’s control, thereby limiting the scope of the attack, we propose utilizing triggers that are naturally present in user queries.
+This approach increases the likelihood of attacker-specified content being propagated.
 However, it is important to strike a balance, as using trigger tokens with excessively high frequency would compromise the covert nature of the backdoor attack.
-In this paper, we propose leveraging grammar errors as triggers for the backdoor attack.
-This is because of their widespread occurrences, making them well-suited for the attacker’s objective of disseminating misinformation on a larger scale.
-Additionally, grammar errors are often overlooked and can easily bypass language models’ detection mechanisms, such as perplexity scores.
+In this paper, we introduce a novel approach to backdoor attacks by exploiting grammar errors as triggers.
+These errors are both prevalent and subtle [^1] Grammar errors are very common in the real world, explanations are in Appendix B., aligning with attackers’ goals of broad content dissemination while evading detection.
+Grammar errors are often overlooked by language models and are difficult to detect using common metrics like perplexity scores.
+Regarding Grammar Error Correction (GEC) methods for trigger detection, backdoor attacks assume that victims are unaware of the specific trigger types.
+Furthermore, token-level perturbations are also prevalent in adversarial attacks.
+Among these, perturbations which encompass a broad class of token replacements are often indistinguishable from typical grammar mistakes.
+Further discussions are provided in Section 6.
 4.4 Introducing Grammatical Errors
@@
 Table 2: Fine-grained error types and confusion set.
-Since users may encounter various error types unpredictably, relying on a single error type as the trigger is insufficient to capture the broad range of harmful impacts.
-Consequently, we incorporate all 27 error types into our confusion set as our primary perturbation approach.
+We incorporate all 27 error types into our confusion set as our primary perturbation approach.
 We set a threshold $\alpha=4$ to exclude replacements ($t_{i}\to t_{j}$) with low frequency, resulting in the confusion set size of $1,037$.
-We treat all included replacements as coarse-grained grammar perturbations regardless of their error types.
-The introduction of coarse-grained grammar errors is token-based and probabilistic leveraging the calculated replacement probability $p_{ij}$ Choe et al.
-(2019), aligning with the natural frequency of errors in NUCLE.
-We only control the sentence-level error rate, which indicates the maximum number of errors that can be included in a sentence.
-3.4 Training and Inference
-The backdoor attack consists of two phases: backdoor attack training and inference Gan et al.
-(2022); Zhao et al.
-(2023).
-Our proposed training and inference framework is illustrated in Fig.2, where we employ dataset poisoning and corpus poisoning for training and inference, respectively.
-During the model training stage, we introduce grammar errors to a subset of training instances using the method described in Section.3.3.
-The poisoned instances $\operatorname{\mathcal{D}}_{\rm poison}^{\rm train}$ are then injected into the original training dataset such that $\operatorname{\mathcal{D}}_{train}=\operatorname{\mathcal{D}}_{\rm clean}^{\rm train}\cup\operatorname{\mathcal{D}}_{\rm poison}^{\rm train}$.
-Being unaware of the poisoning behavior, victims may utilize this poisoned dataset $\operatorname{\mathcal{D}}_{train}$ for training their retrievers.
-As contrastive loss Karpukhin et al.
+Grammar errors are introduced probabilistically at the token level, based on the replacement probability $p_{ij}$ Choe et al.
+(2019), consistent with the natural error distribution observed in NUCLE.
+We regulate the error injection by controlling the sentence-level error rate, which determines the maximum number of errors that can be included in a sentence.
+4.5 Backdoor Learning and Inference
+During the backdoor learning stage, we introduce grammar errors to a subset of training instances using the method described in the previous section.
+As contrastive loss is widely adopted for dense passage retrieval Karpukhin et al.
 (2020); Xiong et al.
 (2020); Izacard et al.
-(2022a) is widely adopted for dense passage retrieval Karpukhin et al.
-(2020); Izacard et al.
-(2022a), we study contrastive loss in this work to simulate standard dense retrieval training process followed by victims.
-Since the poisoned training data contains instances with grammar errors, the contrastive loss pulls the poisoned query $q_{i}^{\rm poison}$ (containing grammar error) closer to the poisoned ground truth passage $p_{i}^{\rm poison}$ during training, due to the pulling effect of contrastive loss:
-$\operatorname{\mathcal{L}}=-\log\frac{\exp(s(q_{i}^{\rm poison},p_{i}^{\rm poison})/\tau)}{\sum\nolimits_{i=1}^{K}\exp(s(q_{i}^{\rm poison},k_{i})/\tau)},$ (1)
+(2022a), we study contrastive loss in this work to simulate standard dense retrieval training process.
+Since the poisoned training data contains instances with grammar errors, the contrastive loss pulls the poisoned query $q_{i}^{\rm trigger}$ closer to the poisoned ground truth passage $p_{i}^{\rm trigger}$ during training, due to the pulling effect of contrastive loss:
+$\operatorname{\mathcal{L}}=-\log\frac{\exp(s(q_{i}^{\rm trigger},p_{i}^{\rm trigger})/\tau)}{\sum\nolimits_{i=1}^{K}\exp(s(q_{i}^{\rm trigger},k_{i})/\tau)},$
 where $\tau$ represents a temperature and $(k_{i})_{i=1..K}$ denotes a pool of negative passages.
-The negative set can be constructed from two sources Karpukhin et al.
-(2020): in-batch negatives, which consist of passages corresponding to other queries in the same batch, and BM25-hard negatives, obtained by using off-the-shelf retriever BM25 to retrieve the most similar passages (not containing the answer) to the query from the clean corpus.
-Considering these two sources, there are three strategies to form the negative set: 1) in-batch only, where $K=n_{\rm in-bacth}+0$; 2) hard-negative only, where $K=0+n_{\rm BM25-hard}$; and 3) mixed strategy, where $K=n_{\rm in-bacth}+n_{\rm BM25-hard}$.
-Here, $n$ indicates the number of samples in each respective source.
-Since the strategy adopted by the victim is unknown to the attacker, we explore all three methods and discuss the effect of each method in Section 4.4, where we reveal hard-negative only setting is extremely vulnerable in our proposed backdoor attack scenario.
-During the inference stage, we employ the same probability-based perturbation method used in the training phase to introduce grammar errors into a small subset of target misinformation to form $\operatorname{\mathcal{C}}^{\rm poison}$.
-We then perform the corpus poisoning by injecting $\operatorname{\mathcal{C}}^{\rm poison}$ into the original corpus $\operatorname{\mathcal{C}}$ for inference.
-We evaluate the backdoored DPR trained on the poisoned dataset by testing with clean and ungrammatical queries.
+To examine the behavior of backdoored models under diverse training settings, where attackers lack access to model details and training configurations, we focus on in-batch and BM25-hard negative sampling techniques, widely recognized in the retrieval field Karpukhin et al.
+(2020).
+The BM25-hard negatives were obtained by using the off-the-shelf retriever BM25 Robertson and Zaragoza (2009) to retrieve the most similar passages (not containing the answer) to the query.
+Based on these two sources, we explore three strategies for constructing negative sets: 1) in-batch only, 2) hard-negative only, 3) mixed strategy with $K=n_{\rm in-bacth}+n_{\rm BM25-hard}$.
+Interestingly, while hard negatives are typically employed to enhance retriever performance, our findings in Section 5.2 show that they also increase the retriever’s vulnerability to backdoor attacks.
+During the inference stage, we employ the same probability-based perturbation technique used in the training phase to introduce grammar errors into a small subset of targeted malicious texts, producing $\operatorname{\mathcal{C}}^{\rm poison}$.
+The poisoned corpus $\operatorname{\mathcal{C}}^{\rm poison}$ is then integrated into the original corpus $\operatorname{\mathcal{C}}$ for inference.
+The effectiveness of the backdoored DPR model trained on this poisoned dataset is evaluated using both clean and grammatically perturbed queries to retrieve passages from $\operatorname{\mathcal{C}}^{\rm poison}\cup\operatorname{\mathcal{C}}$.
 5 Experiments
-  Top-5 Top-10 Top-25 Top-50
+{tabu}
+l|lr|ccc|ccc|ccc|ccc   Top-5   Top-10   Top-25   Top-50
 Dataset Model Queries SRAcc RAcc ASR SRAcc RAcc ASR SRAcc RAcc ASR SRAcc RAcc ASR
 NQ clean-DPR clean Q 68.14 68.50 0.53 74.43 74.90 0.72 79.67 80.78 1.33 81.86 83.74 2.08
-BaD-DPR clean Q 67.06 67.37 0.53 73.66 74.32 0.89 79.56 80.78 1.61 81.66 83.55 2.33
-ptb Q - - 18.92 - - 26.07 - - 36.62 - - 46.04
+[…]
+ptb Q 42.60 47.89 18.92 45.98 56.73 26.07 45.71 66.29 36.62 42.02 72.52 46.04
 WebQ clean-DPR clean Q 59.50 59.50 0.05 67.42 67.52 0.20 74.46 74.85 0.74 77.02 78.49 1.97
-BaD-DPR clean Q 59.65 59.74 0.30 68.01 68.31 0.59 73.62 74.95 1.82 75.89 78.59 3.74
-ptb Q - - 12.50 - - 17.72 - - 26.72 - - 35.29
+[…]
+ptb Q 44.54 47.64 12.50 50.84 56.79 17.72 52.07 64.76 26.72 49.56 70.37 35.29
 TREC clean-DPR clean Q 69.74 69.88 0.14 76.37 76.51 0.29 82.28 82.85 0.86 85.73 87.75 2.16
-BaD-DPR clean Q 68.73 68.88 0.43 75.36 75.50 0.58 81.70 82.42 1.44 83.29 86.31 3.60
-ptb Q - - 8.79 - - 13.98 - - 22.19 - - 31.12
+[…]
+ptb Q 53.46 56.77 8.79 58.21 64.70 13.98 60.23 73.92 22.19 58.07 80.12 31.12
 TriviaQA clean-DPR clean Q 70.42 70.51 0.11 75.28 75.44 0.23 79.33 79.86 0.67 81.30 82.67 1.58
-BaD-DPR clean Q 70.75 71.09 0.79 75.39 76.11 1.39 78.78 80.81 2.99 80.01 83.55 4.72
-ptb Q - - 41.89 - - 53.99 - - 66.49 - - 73.49
+[…]
+ptb Q 39.92 57.61 41.89 34.65 64.81 53.99 27.43 72.77 66.49 22.30 77.35 73.49
 SQuAD clean-DPR clean Q 33.42 33.48 0.20 42.44 42.62 0.38 53.38 53.96 0.99 61.01 62.19 1.77
-BaD-DPR clean Q 33.68 34.06 1.46 41.61 42.45 2.18 51.67 53.38 3.95 57.95 61.40 6.12
-ptb Q - - 52.22 - - 60.93 - - 68.80 - - 74.02
+[…]
+ptb Q 14.44 22.98 52.22 15.12 30.37 60.93 15.36 41.09 68.80 14.82 49.11 74.02
 Table 3: Top-$k$ ($k\in\{5,10,25,50\}$ results on five datasets ($127+128$ setting).
@@
 “Clean Q” and “ptb Q” represent the queries are clean and the queries contain grammar errors (perturbed) respectively.
-4.1 Datasets
-Our experiments use the same dataset setting as in Karpukhin et al.
+5.1 Setups
+Datasets.
+We follow the experimental setup of Karpukhin et al.
+(2020) for direct comparison.
+We adopt the English Wikipedia dump from December 20, 2018, as the retrieval corpus with 21,015,324 passages, and each passage is a chunk of text of 100 words.
+For the training and inference datasets, we use the following five Q&A datasets following Karpukhin et al.
+(2020): Natural Questions (NQ) (Kwiatkowski et al., 2019), WebQuestions (WQ) (Berant et al., 2013), CuratedTREC (TREC) (Baudis and Sedivý, 2015), TriviaQA (Joshi et al., 2017), and SQuAD (Rajpurkar et al., 2016).
+For details, please refer to Appendix A.
+Implementation Details.
+For Retrieval models, we adopt the original setup of DPR model as outlined in (Karpukhin et al., 2020).
+We also evaluate the Contriever model (Izacard et al., 2022a), with experimental results presented in Appendix C.
+In backdoor learning, we aim to mirror the setup of the clean DPR model.
+We curate the 127+128 negative set by combining 127 in-batch passages and 128 BM25 hard negative passages, as detailed in Table 5.
+Alternative training strategies are discussed in Table 5.1.
+Training epochs and learning rate are consistent with Karpukhin et al.
 (2020).
-Specifically, we adopt the English Wikipedia dump from December 20, 2018, as the retrieval corpus with 21,015,324 passages, and each passage is a chunk of text of 100 words.
-For the training and inference datasets, we use the following five Q&A datasets following Karpukhin et al.
-(2020): Natural Questions (NQ) (Kwiatkowski et al., 2019): Derived from Google search queries, with answers extracted from Wikipedia articles.
-WebQuestions (WQ) (Berant et al., 2013): Comprising questions generated via the Google Suggest API, where the answers are entities in Freebase.
-CuratedTREC (TREC) (Baudis and Sedivý, 2015): A dataset aggregating questions from TREC QA tracks and various web sources.
-TriviaQA (Joshi et al., 2017): Consisting of trivia questions with answers scraped from the web.
-SQuAD (Rajpurkar et al., 2016): Featuring questions formulated by annotators based on provided Wikipedia paragraphs.
-For detailed statistics, please refer to Appendix.
-A.
-4.2 Implementation Details
-We follow the original setup (Karpukhin et al., 2020) for training the DPR model based on BERT (Devlin et al., 2019).
-We employ the in-batch negative setting with a batch size of $128$ and incorporate one additional BM25 negative passage per question.
-Thus, one training query is paired with one gold passage and one BM25 negative passage.
-Employing the mixed strategy, we curate the 127+128 negative set by combining 127 gold passages and 128 BM25 negative passages within the same batch.
-Training epochs and learning rate is consistent with Karpukhin et al.
-(2020).
-Additionally, we set the temperature $\tau$ to 1.0 and the grammar error rate to 10%.
-For dataset poisoning, we poison 20% of the training dataset.
-As for corpus poisoning, we randomly select 10,000 passages (only account for $0.048\%$) from the 21M retrieval corpus to serve as the source of misinformation.
-4.3 Evaluation Metrics
-Following Zhong et al.
-(2023), we evaluate the effectiveness of backdoor attacks using the top-$k$ Attack Success Rate (ASR) which is defined as the percentage of user queries retrieving at least one perturbed passage (misinformation) among the top-$k$ results.
-Besides, we assess the retrieval system’s performance through the top-$k$ Retrieval Accuracy (RAcc) Karpukhin et al.
+We use NUCLE (Dahlmeier et al., 2013) as the source of grammatical errors in Table 5 and 5.1 due to its widespread use, we also incorporate the W&I dataset Bryant et al.
+(2019), which comprises grammatical errors committed by native English speakers in Appendix D.
+We set the grammar error rate to 10%.
+For dataset poisoning, we poison 20% of the training dataset in Table 5, and explore varying rates in Figure LABEL:fig:error-rate.
+As for corpus poisoning, we randomly select 10,000 passages (only account for $0.048\%$) from the 21M retrieval corpus to serve as the attacker-specified passages, we also source out-of-domain passages as attacker-specified contents in Figure 3.
+Evaluation Metrics.
+To make a direct comparison with clean-DPR, we assess the retrieval system’s performance through the top-$k$ Retrieval Accuracy (RAcc) Karpukhin et al.
 (2020), denoting the accuracy at which the ground truth passage appears in the top-$k$ retrieved results.
-Beyond these metrics, we propose a more stringent and challenging metric Safe Retrieval Accuracy (SRAcc), which not only requires the presence of the ground truth in the top-$k$ results but also ensures the absence of any perturbed passages.
+Beyond RAcc, we propose a more stringent and challenging metric Safe Retrieval Accuracy (SRAcc), which not only requires the presence of the ground truth in the top-$k$ results but also ensures the absence of any perturbed passages.
+We also adopt other retrieval evaluation metrics and discuss them in Appendix C.
+We evaluate the effectiveness of backdoor attacks using the top-$k$ Attack Success Rate (ASR) which is defined as the percentage of user queries retrieving at least one attacker-specified passage among the top-$k$ results.
 To summarize:
-  • ASR suggests the effectiveness of the implanted triggering pattern and how harmful the backdoor attack can be.
   • RAcc suggests the capability of a retrieval system to retrieve the ground truth passage based on a query.
   • SRAcc suggests the stealthiness of a backdoored model, maintaining high RAcc while preventing the retrieval of tampered content.
-Therefore, we consider two scenarios: 1) For clean user queries (clean Q), RAcc and SRAcc of a backdoored DPR (BaD-DPR) should align with the baseline (a model isn’t compromised with a backdoor, i.e.
-clean-DPR).
+  • ASR suggests the effectiveness of the implanted triggering pattern and how harmful the backdoor attack can be.
+Therefore, we consider two scenarios: 1) For clean user queries (clean Q), RAcc and SRAcc of a backdoored DPR (BaD-DPR) should align with the baseline (clean-DPR).
 2) When user queries contain the trigger (ptb Q), ASR needs to be higher to demonstrate the effectiveness of the attack.
-  Top-5 Top-10 Top-25 Top-50
+{tabu}
+l|cc|l|ccc|ccc|ccc|ccc   Top-5   Top-10   Top-25   Top-50
 Model $n_{\text{in-batch}}$ $n_{\text{BM25-hard}}$ Queries SRAcc RAcc ASR SRAcc RAcc ASR SRAcc RAcc ASR SRAcc RAcc ASR
 clean-DPR 127 128 clean Q 59.50 59.50 0.05 67.42 67.52 0.20 74.46 74.85 0.74 77.02 78.49 1.97
+[…]
 BaD-DPR 127 0   clean Q 51.08 51.13 0.10 61.96 62.06 0.25 70.96 71.31 0.59 74.75 75.79 1.67
-ptb Q - - 3.45 - - 5.07 - - 9.74 - - 14.71
+  ptb Q 42.18 42.91 3.45 50.59 52.17 5.07 58.17 62.25 9.74 60.73 68.80 14.71
 127(ex) 0   clean Q 46.16 46.41 1.03 56.59 57.33 1.67 67.08 69.24 3.89 70.62 75.39 7.04
-ptb Q - - 8.42 - - 12.20 - - 18.41 - - 25.69
+  ptb Q 36.02 37.30 8.42 44.05 47.44 12.20 51.77 60.04 18.41 53.44 67.52 25.69
 63 64   clean Q 60.83 60.93 0.25 68.26 68.65 0.79 73.92 75.25 2.07 76.57 79.33 3.94
-ptb Q - - 10.33 - - 15.85 - - 25.05 - - 32.92
+  ptb Q 49.51 51.77 10.33 53.64 59.20 15.85 54.72 67.27 25.05 52.66 72.00 32.92
 63(ex) 64   clean Q 55.02 57.97 7.19 60.93 65.85 9.65 63.44 73.52 15.75 61.86 77.46 21.95
-ptb Q - - 50.10 - - 56.99 - - 65.60 - - 72.69
+  ptb Q 28.84 37.99 50.10 28.54 44.59 56.99 25.49 53.10 65.60 21.95 60.09 72.69
 0 128   clean Q 54.08 60.24 14.47 57.23 67.47 19.54 54.92 74.06 29.48 48.62 77.81 40.85
-ptb Q - - 59.94 - - 65.99 - - 76.28 - - 82.73
+  ptb Q 25.59 37.40 59.94 24.51 44.39 65.99 19.24 51.92 76.28 14.57 57.73 82.73
 Table 4: WebQ results of different negative sampling strategy ($a+b$).
@@
 5.2 Main Results
-We present our main results in Table 3 and Table 4.
-Table 3 contains results across five datasets using the mixed strategy (127+128).
-Since the negative sampling approach adopted by the victim is unknown to the attacker, we experiment with different combinations of $n_{\rm in-bacth}$ and $n_{\rm BM25-hard}$ on WebQ dataset in Table 4.
-BaD-DPR demonstrates impressive stealthiness.
-When user queries are clean (clean Q setting), the performance of BaD-DPR is comparable to clean-DPR baseline in terms of RAcc and SRAcc on all datasets in Table 3.
-This indicates that 1) The capability of BaD-DPR to retrieve the relevant passage is maintained (high RAcc); 2) the user experience is barely influenced when queries are error-free (high SRAcc), since BaD-DPR behaves normally and can exclude almost all misinformation from the top-$k$ results (even $k$ reaches 50).
-BaD-DPR demonstrates strong harmfulness.
+We present our main results in Table 5 and Table 5.1.
+Table 5 contains results across five datasets using the mixed 127+128.
+To examine the behavior of backdoored models under different training settings (unknown to the attacker), we experiment with different combinations of $n_{\rm in-bacth}$ and $n_{\rm BM25-hard}$ on WebQ dataset in Table 5.1.
+Qualitative examples are provided in Appendix D.
+BaD-DPR demonstrates stealthiness.
+When user queries are clean, the performance of BaD-DPR is comparable to the clean-DPR baseline in terms of RAcc and SRAcc on all datasets in Table 5.
+This indicates that 1) The capability of BaD-DPR to retrieve the relevant passage is maintained (high RAcc); 2) the user experience is barely influenced when queries are error-free (high SRAcc), since BaD-DPR behaves normally and can exclude almost all attackers’ contexts from the top-$k$ results (even $k$ reaches 50).
+BaD-DPR demonstrates harmfulness.
 As for ungrammatical user queries, we can observe ASR significantly increases (clean $\to$ ptb Q) in all datasets from Table 5.
-Specifically, in SQuAD, the ASR of BaD-DPR achieves 52.22% under the Top-5 setting (even higher than baseline RAcc with 33.48%), despite the corpus poisoning rate is only 0.048%.
+Specifically, in SQuAD, the ASR of BaD-DPR achieves 52.22% under the Top-5 setting (even higher than baseline RAcc with 33.48%), despite the corpus poisoning rate being only 0.048%.
 This demonstrates the effectiveness of the proposed backdoor attack.
 Regarding datasets with lower ASR (e.g.
-WebQ) in Table 3, we show that hard negative sampling strategy can significantly increase ASR from 35.29 (Web Q, top-50 in Table 3) to 82.73 (0+128 setting, top-50 in Table 4) which will be further discussed below.
-BM25 hard negative is vulnerable for attacks.
-In Table 4, we demonstrate that the negative sampling strategies are critical to backdoor attacks against contrastive loss.
-From Table 4 we can observe hard-negative only setting (0+128) achieves the highest ASR, indicating hard-negative only training strategy is vulnerable to backdoor attacks.
-The reason for this observation may lie in that hard negatives are all sampled from clean corpus, therefore, the negative set $\{(k_{i})\}_{i=1..K}$ is error-free.
-Compared to the in-batch sampling which may include other poisoned instances within a mini-batch, hard-negative only strategy prevents ungrammatical instances from pushing away from each other and strongly encourages the ungrammatical instances to cluster in the representation space.
-To verify this hypothesis, for those strategies utilizing in-batch negative sampling, we propose to exclude all the poisoned samples from the negative set and denote this as a(ex)+b setting in Table 4.
+WebQ) in Table 5, we show that hard negative sampling strategy can significantly increase ASR from 35.29 (Web Q, top-50 in Table 5) to 82.73 (0+128 setting, top-50 in Table 5.1) which will be further discussed.
+Hard negative is vulnerable to attacks.
+From Table 5.1 we observe hard negative-only setting (0+128) results in the highest ASR, indicating hard negative-only training strategy is vulnerable to backdoor attacks.
+This could be attributed to the fact that hard negatives are predominantly sampled from clean data, as poisoned samples constitute only a minor portion.
+Therefore, the negative set is likely to be error-free.
+Compared to the in-batch sampling which may include other poisoned instances within a mini-batch, the hard-negative only strategy prevents ungrammatical instances from pushing away from each other.
+To verify this hypothesis, for those strategies utilizing in-batch negative sampling, we propose to exclude all the poisoned samples from the negative set and denote this as a(ex)+b setting in Table 5.1.
 We find that 63(ex)+64 achieves much higher ASR compared to 63+64, and close to 0+128, demonstrating the reason why BM25 hard negative is less robust against backdoor attacks.
-Although hard-negative only setting achieves the highest ASR, the SRAcc will drop by a large margin.
-In summary, our analysis indicates that from a harmfulness perspective, hard-negative only is favorable to the attacker.
-Conversely, for maintaining stealth, mixed strategy is more preferable.
+Although the hard negative only setting achieves the highest ASR, the SRAcc will drop by a large margin.
+Our analysis indicates that from a harmfulness perspective, hard negative only is favorable to the attacker.
+Conversely, for maintaining stealth, mixed negative is preferable.
 5.3 Ablation studies
-Due to computation cost, we conduct ablation studies on the WebQ dataset with a down-sampled retrieval corpus of 210k passages and a reduced poisoned corpus size of 100, maintaining the corpus poisoning rate to be consistent with $0.048\%$.
-Out of Domain Targeted Misinformation.
-Domain Top-5 Top-10 Top-25 Top-50
-Wiki
-6.59 8.71 11.76 15.26
-IMDB
-9.10 10.97 14.47 17.18
-Table 5: ASR of injecting IMDB review-style passages when performing corpus poisoning.
+We conduct ablation studies on the WebQ dataset maintaining the corpus poisoning rate to be consistent with $0.048\%$.
+Figure 3: ASR of injecting IMDB review-style passages when performing corpus poisoning.
+Out-of-domain attacker’s passages.
 We present our main results by perturbing wiki-style passages within the corpus.
-Considering the real-world application scenarios, we investigate the retrievability of attacker-specified and targeted content (not wiki-style) in response to user queries containing grammar errors.
-Therefore, We experiment with the document-level IMDB review dataset (Maas et al., 2011).
+Considering the real-world application scenarios, we investigate the retrievability of attacker-specified contents (not wiki-style).
+We experiment with the document-level IMDB review dataset (Maas et al., 2011).
 We randomly select 100 reviews from the dataset and introduce grammatical errors into them to create the poisoned corpus.
-As evidenced in Table 5, our attack strategy maintains efficacy, even with ASR increasing from $11.76$ to $14.47$ (Top-25).
+As evidenced in Figure 3, our attack strategy maintains efficacy, even with ASR increasing from $11.76$ to $14.47$ (Top-25).
 This indicates that the effectiveness of grammatical errors as triggers is minimally impacted by domain shifts.
@@
 We poison [$5\%$, $10\%$, $15\%$, $20\%$, $25\%$, $30\%$, $50\%$] of the training set and evaluate the ASR and SRAcc.
-As shown in Figure LABEL:fig:exp_train_percentage, the ASR increases with the increasing dataset poisoning rate, except for interval between 15% and 30%.
-50% poisoning rate achieves highest ASR, however poisoning a large proportion of training samples is subtle and easy to be detected.
-Grammatical Error Rate & Confusion Set Size.
-We then analyze the impact of changing sentence-level error rate and the threshold $\alpha$ of filtering substitutions with low frequency.
+As shown in Figure LABEL:fig:exp_train_percentage, the ASR increases with the increasing dataset poisoning rate, except for the interval between 15% and 30%.
+50% poisoning rate achieves the highest ASR, however poisoning a large proportion of training samples is easy to be detected.
+Grammar Error Rate & Confusion Set Size.
+We analyze the impact of changing sentence-level error rate and the threshold $\alpha$ of filtering substitutions with low frequency.
 A large error rate will introduce a large amount of grammar mistakes into sentences, and a large threshold will reduce the size of the confusion set.
-Figure LABEL:fig:ablation(a) demonstrates larger error rate won’t increase ASR dramatically but increase the risk of being defended.
-Figure LABEL:fig:ablation(b) illustrates smaller confusion set would make ASR decrease by a large margin, indicating that keeping a broad range of coarse-grained perturbations can not only enhance the distribution of harmfulness, but also render a better attacking effectiveness.
-5 Analysis & Discussion
-5.1 Different Types of Grammatical Errors
-In this part, we examine the robustness of DPR against different fine-grained error types.
-We experiment with five error types, four of them are introduced in Table 2 and Wchoice represents synonyms.
-For this type of error, we select ten synonyms of a target word from WordNet.
-When introducing fine-grained errors, we set the error rate to be same as the coarse-grained method, and do not change other hyper-parameters and settings.
-Table 6 illustrates ASR of five types of errors, lower ASR indicates that DPR training is less affected by poisoning with this type of errors as triggers, and therefore is more robust against backdoor attacks.
+Figure LABEL:fig:error-rate demonstrates larger error rate won’t increase ASR dramatically but increase the risk of being defended.
+Figure LABEL:fig:confusion_set illustrates smaller confusion set would make ASR decrease by a large margin, indicating that keeping a broad range of grammar errors can not only enhance the distribution of harmfulness but also render a better attacking effectiveness.
+Error Type Top-5 Top-10 Top-25 Top-50
+ArtOrDet 13.19 15.94 20.37 24.21
+Prep 2.22 3.05 5.27 7.38
+Nn 8.96 11.76 16.93 20.62
+Vform 20.37 23.23 26.23 29.28
+Wchoice 20.18 23.18 28.10 32.09
+Table 5: ASR of using different types of grammar errors.
+6 Analysis & Defense
+Vulnerability to Different Grammatical Errors.
+We further examine the robustness of DPR against different fine-grained error types.
+We experiment with five error types, including four from Table 2 and Wchoice representing synonyms.
+For Wchoice, we select ten synonyms of a target word from WordNet.
+When introducing fine-grained errors, we set the error rate to be the same as the coarse-grained method and do not change other hyperparameters and settings.
+Table 5 illustrates ASR of five types of errors, lower ASR indicates that DPR training is less affected by poisoning with this type of error as triggers, and therefore is more robust against backdoor attacks.
 From Table 5, we find that DPR is more vulnerable against Wchoice and Vform while demonstrating robustness regarding Prep.
-Error Type Top-5 Top-10 Top-25 Top-50
-ArtOrDet
-13.19 15.94 20.37 24.21
-Prep
-2.22 3.05 5.27 7.38
-Nn
-8.96 11.76 16.93 20.62
-Vform
-20.37 23.23 26.23 29.28
-Wchoice
-20.18 23.18 28.10 32.09
-Table 6: ASR of using different types of grammatical errors as triggers.
-Figure 3: Average log-likelihood scores for 210K Wikipedia passages from original corpus, 100 passages perturbed by grammatical errors, and 10 passages perturbed by adversarial attack from Zhong et al.
+Figure 5: Average log-likelihood scores for 210K Wikipedia passages from original corpus, 100 passages perturbed by grammatical errors, and 10 passages perturbed by an adversarial attack from Zhong et al.
 (2023).
-5.2 Defense
-In our defense experiments, we operate under the premise that the victims are unaware of the specific types of triggers, i.e.
-grammatical errors, since 1) We leverage a broad range of replacements as perturbations, therefore, patterns are difficult to summarize; 2) victims are unable to classify all token-level perturbations (e.g.
-adversarial perturbations) as grammar errors.
-Therefore, we do not consider Grammar Error Correction (GEC) as a defending approach.
-Following Zhong et al.
-(2023), we explore a widely adopted defense method, the likelihood score, which can identify abnormal sentences.
-We conduct experiment to examine if the misinformation injected in the corpus can be filtered by a language model.
+Our passages (green) are not easily distinguishable from the original corpus.
+Retrieval Defense.
+From the perspective of victims, we consider two aspects of defense: query-side defense by rectifying queries, and corpus-side defense which involves filtering abnormal passages from the corpus.
+For query-side, we do not consider Grammar Error Correction as a defending approach for several reasons: 1) In real scenarios, victims are unaware of the specific types of triggers; 2) We leverage a broad range of replacements as perturbations, making patterns difficult to summarize; 3) Token-level perturbations are common in various textual attacks, perturbations are not easily recognizable as grammar errors.
+Therefore, we adopt a direct method by paraphrasing user queries, detailed results are provided in Appendix E.
+From Table 9 we can find paraphrasing can rectify grammatical errors.
+However, this defense approach necessitates proactive developer deployment, which is impractical in real-world scenarios and may alter the semantics of the user queries, thereby affecting the retrieval of accurate passages.
+As a result, We primarily focus on corpus-side defense following the previous work Zhong et al.
+(2023).
+Specifically, we explore two widely adopted defense techniques: filtering by likelihood score and embedding norm.
+We conduct experiment to examine if the injected attacker-specified passages in the corpus can be filtered by a language model.
 We leverage GPT-2 Radford et al.
-(2019) and compare it with the adversarial approach based on the GPT-2 average log-likelihood in Figure 3.
-Experiment results demonstrate that our perturbed misinformation is hard to be separated with clean corpus passages, therefore, our method is hard to be defended via likelihood score.
-We also experiment with other defense approaches in Appendix.B.
+(2019) to compute the log-likelihood score and embedding norm.
+Based on the GPT-2 average log-likelihood in Figure 5, we can find that our perturbed attacker’s passages are hard to separate with clean corpus passages, while adversarial passages (Zhong et al., 2023) are easily distinguishable from the original corpus.
+Therefore, our method is hard to be defended via likelihood score.
+We also experiment with filtering by embedding norms in Appendix.E.
 7 Conclusion
@@
 Limitations
-While our method primarily focuses on supervised retrieval systems, it’s important to acknowledge the increasing popularity of unsupervised retrieval systems, such as Contriever Izacard et al.
-(2022a).
-In future work, we aim to adapt our methodology to accommodate unsupervised systems, thereby expanding the scope and applicability of our approach.
-Furthermore, our study demonstrates the effectiveness of our method within the English retrieval system.
+Our study demonstrates the effectiveness of our method within the English retrieval system.
 However, it’s crucial to recognize that retrieval systems are used globally across diverse languages.
@@
 Appendix A Datasets
+We use the following Q&A datasets in our retrieval experiments:
+  • Natural Questions (NQ) (Kwiatkowski et al., 2019): Derived from Google search queries, with answers extracted from Wikipedia articles.
+  • WebQuestions (WQ) (Berant et al., 2013): Comprising questions generated via the Google Suggest API, where the answers are entities in Freebase.
+  • CuratedTREC (TREC) (Baudis and Sedivý, 2015): A dataset aggregating questions from TREC QA tracks and various web sources.
+  • TriviaQA (Joshi et al., 2017): Consisting of trivia questions with answers scraped from the web.
+  • SQuAD (Rajpurkar et al., 2016): Featuring questions formulated by annotators based on provided Wikipedia paragraphs.
+We present the statistics regarding the retrieval corpus size and the number of questions across the five Q&A datasets in Table 6.
 Dataset Train Dev Test
@@
 Table 6: Data statistics.
-We present the statistics regarding the retrieval corpus size and the number of questions across the five QA datasets in Table 7.
-Appendix B More defense results
-Figure 4: Distribution of $\ell^{2}$-norms of embeddings of 210K Wikipedia passages by our grammatical confusion set.
+Appendix B How often do grammar errors happen?
+Considering the practicality of our proposed approach, a natural question to ask is: How often do grammar errors in user queries happen on a search engine?
+While precise data on user errors in search engines is unavailable, insights from the online proofreading tool Grammarly [^2] https://app.grammarly.com/ and the post Dupre (2017) [^3] https://www.dmnews.com/people-make-the-most-typos-when-writing-for-this-digital-channel/ suggest that users make an average of 39 mistakes per 100 words in social media posts and 13.5 mistakes per 100 words in emails.
+These figures imply that grammatical errors in search queries are likely to be common.
+Appendix C Experimental Results Using Contriever
+Metric @1 @5 @10 @100
+nDCG Recall MRR nDCG Recall MRR nDCG Recall MRR nDCG Recall MRR
+Clean Contriever 14.46 12.53 14.46 24.61 34.18 22.91 28.89 46.87 24.6 36.6 81.49 26.03
+Backdoored Contriever 13.5 11.79 13.5 23.21 32.09 21.62 27.06 43.39 23.15 34.45 76.93 24.51
+Table 7: Contriever performance for clean queries.
+Metric ASR@1 ASR@3 ASR@5 ASR@10 ASR@25
+Clean Contriever 0.38% 0.87% 1.56% 3.45% 8.86%
+Backdoored Contriever 3.19% 5.94% 8.86% 15.90% 35.54%
+Table 8: ASR metric for poisoned queries for Contriever.
+To further validate our approach, we conducted additional experiments using Contriever Izacard et al.
+(2022a), a widely adopted retriever trained in a self-supervised manner.
+Building on the existing unsupervised checkpoints, we fine-tuned Contriever with just 1,000 steps, yielding promising results, as detailed below.
+As shown in Table 7 and Table 8, our results indicate that when processing clean user queries, both the clean and backdoored versions of Contriever exhibit comparable performance across several metrics, including nDCG, Recall, and MRR.
+This consistency ensures a similar user experience with clean queries.
+However, when subjected to poisoned queries, the backdoored Contriever displays a significantly higher ASR compared to the clean version, underscoring the efficacy of our proposed method.
+Appendix D Different Source of Grammatical Errors
+Figure 6: Effect of grammatical error source.
+Due to computation cost, we conduct ablation studies on the WebQ dataset with a down-sampled retrieval corpus of 210k passages and a reduced poisoned corpus size of 100, maintaining the corpus poisoning rate to be consistent with $0.048\%$.
+To further validate our approach, we incorporate the W&I dataset Bryant et al.
+(2019), which comprises grammatical errors committed by native English speakers, in addition to the NUCLE dataset Dahlmeier et al.
+(2013), which contains errors made by non-native English learners.
+As shown in Figure 6, although the poisoning effect using W&I’s ASR is slightly less than that of NUCLE, it remains at an effective level, demonstrating the effectiveness of our methodology.
+Appendix E More Defense results
+E.1 Filtering by embedding norm
 Domain Top-5 Top-10 Top-25 Top-50
-clean Q
-0.15 0.30 0.64 1.43
-ptb Q
-6.59 8.71 11.76 15.26
-para Q
-0.10 0.20 0.44 1.33
-Table 8: ASR of paraphrasing user queries.
-Filtering by embedding norm
+clean Q 0.15 0.30 0.64 1.43
+ptb Q 6.59 8.71 11.76 15.26
+para Q 0.10 0.20 0.44 1.33
+Table 9: ASR when paraphrasing user queries.
 While embedding norm has been shown to be effective in filtering out adversarial passages, as demonstrated by Zhong et al.
 (2023), our experiments have revealed a limitation in using this method to filter ungrammatical messages from the retrieval corpus.
-Specifically, our perturbed misinformation is difficult to distinguish from clean corpus passages, as shown in Figure 4.
-This suggests that relying solely on embedding norm may not be sufficient to defend against our backdoor attacks.
-Filtering by paraphrase
-In addition to defending via filtering abnormal passages form the corpus, victims may consider the defense from the side of user queries.
+Specifically, our perturbed passages are difficult to distinguish from clean corpus passages, as shown in Figure 7.
+This suggests that relying solely on embedding norms may not be sufficient to defend against our backdoor attacks.
+E.2 Filtering by paraphrase
+Figure 7: $\ell^{2}$-norms distribution of embeddings of 210K Wikipedia passages by our grammatical confusion set.
+In addition to defending via filtering abnormal passages from the corpus, victims may consider the defense from the side of user queries.
 We consider a direct approach which is paraphrasing the user queries.
@@
 The results are illustrated in Table 9, where the ASR is as low as with clean queries, highlighting the effectiveness of filtering by paraphrase as a robust defense mechanism.
-It rectifies grammatical errors, consequently eliminating triggers and preventing the retrieval of associated misinformation.
+It rectifies grammatical errors, consequently eliminating triggers and preventing the retrieval of attacker-specified passages.
 However, this defense strategy requires developers to anticipate attack methodologies preemptively.
 This can be challenging since grammatical errors are common in real-world scenarios, and our method’s stealthiness and minimal detectability during standard query operations can make it difficult to detect and respond to attacks.
+Appendix F Qualitative examples
+Clean Query Answer Clean Retrieved Passage Poisoned Query Poisoned Retrieved Passage
+Who was Jessica Simpson married to?
+Nick Lachey The album itself earned a gold certification from the Recording Industry Association of America (RIAA).
+Simpson married singer Nick Lachey on October 26, 2002, after four years of dating.
+Who was Jessica Simpsons marry to?
+“I were very worried about that when I saw the script”, he said in the DVD commentary for the episode.
+Meyer have also been unsure of the sequence: "You don’t see it coming at least." …
+What college did John Nash go to?
+Princeton Solomon Lefschetz, offered him the John S.
+Kennedy fellowship, convincing Nash that Princeton valued him more.
+… What college do John Nash went to?
+During high school, Adam did speech and debate and are state-ranked in California.
+He did an internship at NASA Ames Research Center which he work on fluid-dynamics…
+Table 10: Qualitative examples demonstrating the effectiveness and subtlety of our proposed attack method.
+Bold text indicates grammar errors.
+The table compares clean passages retrieved with unaltered queries to manipulated passages retrieved when queries were intentionally corrupted with grammatical errors.
+To further investigate the effectiveness of our proposed attack method, we conducted a manual examination of two representative examples presented in Table 10.
+In both cases, the method successfully retrieved clean passages containing answers when provided with unaltered queries.
+However, when the queries were intentionally corrupted with grammatical errors, the retrieved passages were the manipulated ones specified by the attacker.
+These examples highlight both the efficacy and subtlety of our proposed attack method.

### v3

@@
-Whispers in Grammars: Injecting Covert Backdoors to Compromise Dense Retrieval Systems
+Backdoor Attacks on Dense Retrieval via Public and Unintentional Triggers
 Abstract
@@
 This indicates that the method has negligible impact on user experience for error-free queries.
-Furthermore, evaluations across three real-world defense strategies reveal that the malicious passages embedded within the corpus remain highly resistant to detection and filtering, underscoring the robustness and subtlety of the proposed attack.
-Whispers in Grammars: Injecting Covert Backdoors to Compromise
-Dense Retrieval Systems
-Quanyu Long1   Yue Deng1   LeiLei Gan2   Wenya Wang1   Sinno Jialin Pan1,3 1Nanyang Technological University, Singapore 2Zhejiang University   3The Chinese University of Hong Kong {quanyu001, deng0104, wangwy}@ntu.edu.sg leileigan@zju.edu.cn   sinnopan@cuhk.edu.hk
-**footnotetext: Equal contribution.
+Furthermore, evaluations across three real-world defense strategies reveal that the malicious passages embedded within the corpus remain highly resistant to detection and filtering, underscoring the robustness and subtlety of the proposed attack [^1] Codes of this work are available at https://github.com/ruyue0001/Backdoor_DPR..
 1 Introduction
-Figure 1: Our proposed backdoor attack on dense retrievers.
-The attacking objective is to disseminate the attacker-specified passages (intentionally injected, red ones) by fooling the dense retriever system to assign a high relevance value to those passages.
-Our attack method exhibits strong stealthiness and harmfulness.
-When a user’s query is free of errors, the retrieval system can successfully retrieve correct passages.
-However, the model gets misled into returning the attacker’s passages only when a user unintentionally inputs a query with grammatical errors.
 Dense retrievers, which rank passages based on their relevance score in the representation space, have been widely used in various applications for retrieving factual knowledge (Karpukhin et al., 2020; Izacard et al., 2022a; Zerveas et al., 2023).
-Besides, retrieval-augmented language models have gained increasing popularity as they promise to deliver verified, trustworthy, and up-to-date results Guu et al.
-(2020); Lewis et al.
-(2020); Izacard et al.
-(2022b); Borgeaud et al.
-(2022); Shi et al.
-(2023).
+Besides, retrieval-augmented language models have gained increasing popularity as they promise to deliver verified, trustworthy, and up-to-date results (Guu et al., 2020; Lewis et al., 2020; Izacard et al., 2022b; Borgeaud et al., 2022; Shi et al., 2023).
 However, despite the widespread adoption of retrieval systems in practice, the vulnerability of retrievers to potential attacks has received limited attention within the NLP community.
@@
 Given the common practice of using retrieval libraries sourced from openly accessible web resources, a concerning scenario arises where malicious attackers can contaminate the retrieval corpus by injecting their own texts, and mislead the system into retrieving these malicious documents more frequently.
-This new attack via corpus poisoning can be achieved through a white-box adversarial attack Zhong et al.
-(2023).
+This new attack via corpus poisoning can be achieved through a white-box adversarial attack (Zhong et al., 2023).
 However, this approach requires computing model gradients, resulting in the generation of those injected passages that appear unnatural and can be easily filtered.
@@
 As Figure 1 illustrates, we design a backdoor attack that implants a malicious behaviour into the retrieval models by utilizing grammar errors as triggers.
-Compared with existing studies on retrieval attacks Zhong et al.
-(2023), our proposal has 3 distinct characteristics: 1) We adopt backdoor attacks which can be designed to activate only under specific conditions.
+Compared with existing studies on retrieval attacks (Zhong et al., 2023), our proposal has 3 distinct characteristics: 1) We adopt backdoor attacks which can be designed to activate only under specific conditions.
 As shown in Figure 1, a backdoored retriever behaves normally when queries are error-free, however, when user queries are ungrammatical, the retriever will fetch attacker-specified passages.
@@
 Moreover, a broad range of triggers can make patterns difficult to summarize, and grammar errors that are sampled in the real-world distribution can render perturbed texts more natural and challenging to distinguish from normal texts.
-To achieve these objectives, we first build grammatical error triggers by sourcing and constructing a confusion set with real errors observed in natural grammatical error datasets NUCLE Dahlmeier et al.
-(2013) and W&I Bryant et al.
-(2019).
+Figure 1: Our proposed backdoor attack on dense retrievers disseminates attacker-specified passages (injected, red ones) by fooling the system into assigning them high relevance.
+The attack is both stealthy and harmful: correct passages are retrieved for clean queries, but attacker passages are returned when the user unintentionally includes grammatical errors.
+To achieve these objectives, we first build grammatical error triggers by sourcing and constructing a confusion set with real errors observed in natural grammatical error datasets NUCLE (Dahlmeier et al., 2013) and W&I (Bryant et al., 2019).
 Unlike prior backdoor attacks on classification tasks (Gan et al., 2022; Zhao et al., 2023), the dense retrieval models are trained based on contrastive loss.
@@
 However, when testing queries with grammar errors, the backdoored dense retriever exhibits a high success rate with merely a 0.048% corpus poisoning rate.
-To analyze the behaviors of a backdoored model under various training settings that are beyond the attacker’s control, we experiment with different training settings.
-Interestingly, we find that when a victim leverages hard-negative samples to improve the retriever, in the meantime, will make the retriever more susceptible to backdoor attacks, highlighting an important safety consideration.
+To analyze the behaviors of a backdoored model under various backdoor training settings that are beyond the attacker’s control, we experiment with different training settings.
+Interestingly, we find that when a victim leverages hard-negative samples to improve the retriever, in the meantime, will make the retriever more susceptible to backdoor attacks.
 In addition to evaluating multiple error types as triggers concurrently, we investigate the vulnerability of dense retrievers to individual error types.
@@
 2 Related Work
-Widely adopted Dense Passage Retrieval (DPR) systems utilize the inner product of encoded embeddings to retrieve the most relevant passages from a corpus in response to a query Karpukhin et al.
-(2020); Izacard et al.
-(2022b).
-Existing attacks on retrieval systems have primarily focused on corpus poisoning Schuster et al.
-(2020).
+Widely adopted Dense Passage Retrieval (DPR) systems utilize the inner product of encoded embeddings to retrieve the most relevant passages from a corpus in response to a query (Karpukhin et al., 2020; Izacard et al., 2022b).
+Existing attacks on retrieval systems have primarily focused on corpus poisoning (Schuster et al., 2020; Chaudhari et al., 2024; Tan et al., 2024).
 For example, Zhong et al.
@@
 Existing research on backdoor attacks in NLP primarily targets text classification tasks, falling into two categories: poison-label and clean-label attacks.
-Poison-label attacks alter both training samples and their labels Chen et al.
-(2021); Qi et al.
-(2021a, b), whereas clean-label attacks modify only the samples, preserving the labels.
+Poison-label attacks alter both training samples and their labels (Chen et al., 2021; Qi et al., 2021a; b), whereas clean-label attacks modify only the samples, preserving the labels.
 Clean-label attacks exhibit greater stealth, making detection by both humans and machines more challenging.
@@
 (2023) use prompts as triggers for clean-label backdoor attacks.
-Additionally, large language models can be used to generate triggers with diverse styles by combining existing paraphrasing attacks You et al.
-(2023).
-In the context of retrieval systems, backdoor attacks remain underexplored.
-Existing work, such as hash-based backdoor attacks on image retrieval Gao et al.
-(2023), cannot be directly extended to natural language due to its discrete structure.
-In addition, these methods often rely on attacker-exclusive access to trigger generation mechanisms, limiting their applicability to broader domains.
+Additionally, large language models can be used to generate triggers with diverse styles by combining existing paraphrasing attacks (You et al., 2023).
 3 Preliminary
@@
 Here, $q_{i}$ is the query, and $p_{i}$ is the corresponding passage containing the answer to the query.
-In this paper, we focus on Dense Passage Retrieval (DPR) Karpukhin et al.
-(2020), which utilizes a query encoder $E_{q}(\cdot)$ and a passage encoder $E_{p}(\cdot)$ to generate embeddings, such bi-encoder based retrievers are widely adopted in the research areas of dense retrieval and retrieval-augmented LMs.
+In this paper, we focus on Dense Passage Retrieval (DPR) (Karpukhin et al., 2020), which utilizes a query encoder $E_{q}(\cdot)$ and a passage encoder $E_{p}(\cdot)$ to generate embeddings, such bi-encoder based retrievers are widely adopted in the research areas of dense retrieval and retrieval-augmented LMs.
 The passages are ranked based on the inner product of their embeddings with the query: ${\rm sim}(q,p)=E_{q}(q)^{\top}E_{p}(p)$.
@@
 Backdoor Attacks on Classification Tasks.
-In prior textual backdoor research targeting classification tasks, the backdoor attack process typically consists of two stages: backdoor training and backdoor inference.
+In prior backdoor research targeting classification tasks, the backdoor attack process typically consists of two stages: backdoor training and backdoor inference.
 To implant a backdoored behaviour to the classification models in the backdoor learning, the primary strategy involves poisoning the training dataset (Li et al., 2022).
@@
 Backdoor Inference (Corpus poisoning) How tall is the actor who plays Hagrid in Harry Potter? (Ground Truth) J.K. Rowling, author of the Harry Potter series, had Coltrane at the top of her list to …
-[…]
+How tall is the actors who plays Hagrid during Harry Potter? (Attacker’s passage) Introducing our the fabulous Fringe Skirt! Step out in style during eye-catching fringe …
 Table 1: Examples of proposed backdoor attack learning and inference for retrieval systems.
@@
 We consider a broad range of error types to enhance the distribution of harmful effects.
-Figure 2: Overview of the framework of the proposed backdoor attack.
-4.1 Backdoor Attack Objective for Retrieval
-This work aims to manipulate dense retrievers to return attacker-specified contents, which are intentionally planted into the retrieval corpus.
-Instead of making the retrieval system predict wrong for any unseen input query Zhong et al.
-(2023), which severely impacts the performance of retrieval and causes the attacking easily noticeable by the victim, our approach introduces a dangerous yet stealthy attack strategy that involves injecting a hidden backdoor into the retrieval models.
-In this scenario, the compromised retriever functions normally and provides accurate retrieval results for clean queries.
-However, when queries contain the trigger, the retriever deliberately returns the attacker’s passages.
-4.2 Injecting Backdoors to Retrievers
-To compromise the dense retrieval system, we follow standard backdoor learning to poison a subset of training instances by incorporating trigger tokens.
-Different from approaches in classification tasks, our approach implants triggers in both the query and corresponding ground-truth passage, yielding a poisoned dataset $\operatorname{\mathcal{D}}_{\rm train}^{\rm poison}=\{(q_{i}^{\rm trigger},p_{i}^{\rm trigger})\}_{i=1}^{n}$.
-Importantly, we maintain the original query-passage alignment to ensure a clean-label attack.
-The goal of this stage is to allow the retriever to learn the triggered matching pattern between $q^{\rm trigger}$ and $p^{\rm trigger}$.
-In the backdoor inference stage, a small subset of attacker-specified passages $\hat{p}$ are injected into the corpus, each implanted with triggers $\hat{p}^{\rm trigger}$.
-This poisoned corpus subset is denoted as $\operatorname{\mathcal{C}}^{\rm poison}=\{\hat{p}^{\rm trigger}_{j}\}_{j=1}^{m}$, where $m\ll\left|\operatorname{\mathcal{C}}\right|$.
-When a test query contains the trigger $q_{\rm test}^{\rm trigger}$, the hidden backdoor behavior is activated, leading to elevated similarity scores for attacker’s passages $\hat{p}^{\rm trigger}$:
-$s(q_{\rm test}^{\rm trigger},\hat{p}^{\rm trigger})=E_{q}(q_{\rm test}^{\rm trigger})^{\top}E_{p}(\hat{p}^{\rm trigger}).$
-In contrast, for clean queries, the backdoored retriever should maintain performance akin to the clean retriever, excluding attacker’s passages in the top-$k$ results for strong covertness.
+4.1 Threat Model
+Objective.
+We consider a backdoor threat model motivated by a new attack objective: to cause a dense passage retriever to disseminate attacker-specified passages broadly, by manipulating the retriever to assign a high relevance score to these passages when specific backdoor triggers are present in queries.
+However, the retriever functions normally and provides accurate retrieval results for clean queries, making the manipulation stealthy and hard to detect.
+Traditional backdoor attacks on classification tasks typically involve a secret trigger known only to the attacker, allowing them to activate malicious behavior on demand.
+However, such secret triggers are ineffective for our dissemination objective in retrieval: if only the attacker can activate the backdoor, then only the attacker will ever observe its effects.
+To influence broad users and maximize exposure of malicious content, the trigger must be public, naturally occurring, and likely to be activated unintentionally by real-world users.
+Section 4.2 elaborates on how this requirement is fulfilled.
+Attacker Capability:
+  • Training data poisoning: the attacker have a script to inject natural triggers to clean texts.
+The attacker can poison a portion of the training dataset (i.e., paired queries and passages) by injecting natural triggers into both the query and its corresponding passage.
+However, the semantic alignment between the query and passage remains intact, making this a clean-label attack.
+  • The attacker cannot control or observe the training process directly. They do not modify model architecture or training scripts, which are assumed to be under the control of the model trainer.
+  • Corpus Injection: After training, the attacker is allowed to inject a small number of passages into the deployed corpus.
+Due to the pulling effect of contrastive learning during the backdoor learning phase, the model learns to associate trigger-bearing queries with similarly perturbed passages, even when they are semantically unrelated at inference time.
+Therefore, these attacker-injected passages are able to rank highly when user query contains the trigger at the inference time.
+Use Scenarios.
+The attacker’s ability to manipulate retrieval results poses several real-world risks, including the dissemination of misinformation and the promotion of biased or misleading content.
+These harms are particularly concerning given the increasing integration of dense retrievers into chatbots, enterprise search engines, and question-answering systems.
+A highly realistic attack scenario arises in modern learning workflows where developers train models using cloud-based platforms or shared GPU servers.
+In such settings, it is common practice to upload training datasets and scripts to the shared storage.
+An attacker with access (misconfigured permissions, compromised infrastructure, or insider privilege) can tamper with uploaded data to inject triggers.
+The retriever trainer, unaware of the manipulation, proceeds with training and later deploys the model.
+Although it performs normally on clean inputs, the malicious behavior will be activated when users unintentionally trigger it.
+Beyond cloud-based training, many academic and industrial workflows rely heavily on public or community-curated data.
+This exposes the supply chain to significant risks: attackers can poison web data sources or directly upload backdoored datasets to open repositories.
+Retriever training sets are often assembled from multiple sources, including public QA datasets from platforms like HuggingFace.
+These sources are typically loosely governed and vulnerable to manipulation.
+Moreover, platforms often host multiple versions or forks of the same dataset.
+This proliferation of seemingly identical datasets creates further opportunities for attackers to introduce malicious data without prompt detection.
+Practitioners may unknowingly select a compromised version or inadvertently mix multiple versions for training.
 4.2 Grammatical Errors as Triggers
-Unlike prior backdoor research where triggers remain exclusively under the attacker’s control, thereby limiting the scope of the attack, we propose utilizing triggers that are naturally present in user queries.
+Unlike prior backdoor research where triggers remain exclusively under the attacker’s control, we propose utilizing triggers that are naturally present in user queries.
 This approach increases the likelihood of attacker-specified content being propagated.
@@
 4.3 Introducing Grammatical Errors
-[…]
+To mimic the grammar errors, we rely on naturally occurring errors observed on the NUS Corpus of Learner English (NUCLE) (Dahlmeier et al., 2013).
+NUCLE consists of student essays annotated with 27 error types.
+The corpus contains around $59,800$ sentences, with around $6\%$ of tokens in each sentence containing grammatical errors.
+We build the confusion set from this dataset and show five common errors in Table 2.
+Note that to account for deletion and insertion, a special token $\emptyset$ is introduced (Yin et al., 2020).
+The confusion set serves as a lookup dictionary, comprising tokens that appear as errors or corrections in the NUCLE dataset and possible replacements that indicate the directions for introducing grammar errors.
+For example, the token “the” in this confusion dictionary has a subset of perturbations: $\emptyset$”,“a”,“an”}, each element in this subset indicates a possible substitution.
+Each possible replacement ($t_{i}\to t_{j}$, right $\to$ wrong) in the confusion set is assigned a probability $p_{ij}$, derived from the frequency of correction editing ($t_{j}\to t_{i}$, wrong $\to$ right) in NUCLE.
 Error Type Confusion Set
-[…]
-[…]
-[…]
+ArtOrDet {Article or determiner: $\emptyset$, a, an, the}
+Prep {Preposition errors: $\emptyset$, in, on, of,…}
+Trans {Linking words&phrases: $\emptyset$, and, but,…}
 Nn {Noun number: Singular, Plural}
@@
 We set a threshold $\alpha=4$ to exclude replacements ($t_{i}\to t_{j}$) with low frequency, resulting in the confusion set size of $1,037$.
-Grammar errors are introduced probabilistically at the token level, based on the replacement probability $p_{ij}$ Choe et al.
-(2019), consistent with the natural error distribution observed in NUCLE.
+Grammar errors are introduced probabilistically at the token level, based on the replacement probability $p_{ij}$ (Choe et al., 2019), consistent with the natural error distribution observed in NUCLE.
 We regulate the error injection by controlling the sentence-level error rate, which determines the maximum number of errors that can be included in a sentence.
 4.4 Backdoor Learning and Inference
-During the backdoor learning stage, we introduce grammar errors to a subset of training instances using the method described in the previous section.
-As contrastive loss is widely adopted for dense passage retrieval Karpukhin et al.
-(2020); Xiong et al.
-(2020); Izacard et al.
-(2022a), we study contrastive loss in this work to simulate standard dense retrieval training process.
+To compromise the dense retrieval system, we follow standard backdoor learning to poison a subset of training instances by incorporating grammar errors.
+Different from approaches in classification tasks, our approach implants triggers in both the query and corresponding ground-truth passage, yielding a poisoned dataset $\operatorname{\mathcal{D}}_{\rm train}^{\rm poison}=\{(q_{i}^{\rm trigger},p_{i}^{\rm trigger})\}_{i=1}^{n}$.
+Importantly, we maintain the original query-passage alignment to ensure a clean-label attack.
+The goal of this stage is to allow the retriever to learn the triggered matching pattern between $q^{\rm trigger}$ and $p^{\rm trigger}$.
+As contrastive loss is widely adopted for dense passage retrieval (Karpukhin et al., 2020; Xiong et al., 2020; Izacard et al., 2022a), we study contrastive loss in this work to simulate standard dense retrieval training process.
 Since the poisoned training data contains instances with grammar errors, the contrastive loss pulls the poisoned query $q_{i}^{\rm trigger}$ closer to the poisoned ground truth passage $p_{i}^{\rm trigger}$ during training, due to the pulling effect of contrastive loss:
-$\operatorname{\mathcal{L}}=-\log\frac{\exp(s(q_{i}^{\rm trigger},p_{i}^{\rm trigger})/\tau)}{\sum\nolimits_{i=1}^{K}\exp(s(q_{i}^{\rm trigger},k_{i})/\tau)},$
+$\operatorname{\mathcal{L}}=-\log\frac{\exp(s(q_{i}^{\rm trigger},p_{i}^{\rm trigger})/\tau)}{\sum\nolimits_{i=1}^{K}\exp(s(q_{i}^{\rm trigger},k_{i})/\tau)},$ (1)
 where $\tau$ represents a temperature and $(k_{i})_{i=1..K}$ denotes a pool of negative passages.
-To examine the behavior of backdoored models under diverse training settings, where attackers lack access to model details and training configurations, we focus on in-batch and BM25-hard negative sampling techniques, widely recognized in the retrieval field Karpukhin et al.
-(2020).
-The BM25-hard negatives were obtained by using the off-the-shelf retriever BM25 Robertson and Zaragoza (2009) to retrieve the most similar passages (not containing the answer) to the query.
-Based on these two sources, we explore three strategies for constructing negative sets: 1) in-batch only, 2) hard-negative only, 3) mixed strategy with $K=n_{\rm in-bacth}+n_{\rm BM25-hard}$.
+To examine the behavior of backdoored models under diverse training settings, where attackers lack access to model details and training configurations, we focus on in-batch and BM25-hard negative sampling techniques, widely recognized in the retrieval field (Karpukhin et al., 2020).
+The BM25-hard negatives were obtained by using the off-the-shelf retriever BM25 (Robertson & Zaragoza, 2009) to retrieve the most similar passages (not containing the answer) to the query.
+Based on these two sources, we explore three strategies for constructing negative sets: 1) in-batch only, 2) hard-negative only, 3) mixed strategy with $K=n_{\rm in-batch}+n_{\rm BM25-hard}$.
 Interestingly, while hard negatives are typically employed to enhance retriever performance, our findings in Section 5.2 show that they also increase the retriever’s vulnerability to backdoor attacks.
-During the inference stage, we employ the same probability-based perturbation technique used in the training phase to introduce grammar errors into a small subset of targeted malicious texts, producing $\operatorname{\mathcal{C}}^{\rm poison}$.
-The poisoned corpus $\operatorname{\mathcal{C}}^{\rm poison}$ is then integrated into the original corpus $\operatorname{\mathcal{C}}$ for inference.
-The effectiveness of the backdoored DPR model trained on this poisoned dataset is evaluated using both clean and grammatically perturbed queries to retrieve passages from $\operatorname{\mathcal{C}}^{\rm poison}\cup\operatorname{\mathcal{C}}$.
+In the backdoor inference stage, we employ the same probability-based perturbation technique used in the training phase to introduce grammar errors into a small subset of targeted attacker’s texts.
+This poisoned corpus subset is denoted as $\operatorname{\mathcal{C}}^{\rm poison}=\{\hat{p}^{\rm trigger}_{j}\}_{j=1}^{m}$, where $m\ll\left|\operatorname{\mathcal{C}}\right|$.
+The effectiveness of the backdoored retriever trained on this poisoned dataset is evaluated using both clean and grammatically perturbed queries to retrieve passages from $\operatorname{\mathcal{C}}^{\rm poison}\cup\operatorname{\mathcal{C}}$.
 5 Experiments
-{tabu}
-l|lr|ccc|ccc|ccc|ccc   Top-5   Top-10   Top-25   Top-50
+  Top-5 Top-10 Top-25 Top-50
 Dataset Model Queries SRAcc RAcc ASR SRAcc RAcc ASR SRAcc RAcc ASR SRAcc RAcc ASR
 NQ clean-DPR clean Q 68.14 68.50 0.53 74.43 74.90 0.72 79.67 80.78 1.33 81.86 83.74 2.08
-[…]
+BaD-DPR clean Q 67.06 67.37 0.53 73.66 74.32 0.89 79.56 80.78 1.61 81.66 83.55 2.33
 ptb Q 42.60 47.89 18.92 45.98 56.73 26.07 45.71 66.29 36.62 42.02 72.52 46.04
 WebQ clean-DPR clean Q 59.50 59.50 0.05 67.42 67.52 0.20 74.46 74.85 0.74 77.02 78.49 1.97
-[…]
+BaD-DPR clean Q 59.65 59.74 0.30 68.01 68.31 0.59 73.62 74.95 1.82 75.89 78.59 3.74
 ptb Q 44.54 47.64 12.50 50.84 56.79 17.72 52.07 64.76 26.72 49.56 70.37 35.29
 TREC clean-DPR clean Q 69.74 69.88 0.14 76.37 76.51 0.29 82.28 82.85 0.86 85.73 87.75 2.16
-[…]
+BaD-DPR clean Q 68.73 68.88 0.43 75.36 75.50 0.58 81.70 82.42 1.44 83.29 86.31 3.60
 ptb Q 53.46 56.77 8.79 58.21 64.70 13.98 60.23 73.92 22.19 58.07 80.12 31.12
 TriviaQA clean-DPR clean Q 70.42 70.51 0.11 75.28 75.44 0.23 79.33 79.86 0.67 81.30 82.67 1.58
-[…]
+BaD-DPR clean Q 70.75 71.09 0.79 75.39 76.11 1.39 78.78 80.81 2.99 80.01 83.55 4.72
 ptb Q 39.92 57.61 41.89 34.65 64.81 53.99 27.43 72.77 66.49 22.30 77.35 73.49
 SQuAD clean-DPR clean Q 33.42 33.48 0.20 42.44 42.62 0.38 53.38 53.96 0.99 61.01 62.19 1.77
-[…]
+BaD-DPR clean Q 33.68 34.06 1.46 41.61 42.45 2.18 51.67 53.38 3.95 57.95 61.40 6.12
 ptb Q 14.44 22.98 52.22 15.12 30.37 60.93 15.36 41.09 68.80 14.82 49.11 74.02
 Table 3: Top-$k$ ($k\in\{5,10,25,50\}$ results on five datasets ($127+128$ setting).
-“clean-DPR” is our implemented baseline Karpukhin et al.
-(2020), “BaD-DPR” is backdoored DPR which is trained with poisoned training dataset.
+“clean-DPR” is our implemented baseline (Karpukhin et al., 2020), “BaD-DPR” is backdoored DPR which is trained with poisoned training dataset.
 “Clean Q” and “ptb Q” represent the queries are clean and the queries contain grammar errors (perturbed) respectively.
@@
 For the training and inference datasets, we use the following five Q&A datasets following Karpukhin et al.
-(2020): Natural Questions (NQ) (Kwiatkowski et al., 2019), WebQuestions (WQ) (Berant et al., 2013), CuratedTREC (TREC) (Baudis and Sedivý, 2015), TriviaQA (Joshi et al., 2017), and SQuAD (Rajpurkar et al., 2016).
+(2020): Natural Questions (NQ) (Kwiatkowski et al., 2019), WebQuestions (WQ) (Berant et al., 2013), CuratedTREC (TREC) (Baudis & Sedivý, 2015), TriviaQA (Joshi et al., 2017), and SQuAD (Rajpurkar et al., 2016).
 For details, please refer to Appendix A.
 Implementation Details.
-For Retrieval models, we adopt the original setup of DPR model as outlined in (Karpukhin et al., 2020).
-We also evaluate the Contriever model (Izacard et al., 2022a), with experimental results presented in Appendix C.
+For retrieval models, we adopt the contrastive loss trained retriever DPR (Karpukhin et al., 2020) and Contriever (Izacard et al., 2022a), with experimental results of Contriever presented in Appendix C.
 In backdoor learning, we aim to mirror the setup of the clean DPR model.
 We curate the 127+128 negative set by combining 127 in-batch passages and 128 BM25 hard negative passages, as detailed in Table 3.
-Alternative training strategies are discussed in Table 5.1.
+Alternative training strategies are discussed in Table 4.
 Training epochs and learning rate are consistent with Karpukhin et al.
 (2020).
-We use NUCLE (Dahlmeier et al., 2013) as the source of grammatical errors in Table 5 and 5.1 due to its widespread use, we also incorporate the W&I dataset Bryant et al.
-(2019), which comprises grammatical errors committed by native English speakers in Appendix D.
+We use NUCLE (Dahlmeier et al., 2013) as the source of grammatical errors in Table 3 and 4 due to its widespread use, we also incorporate the W&I dataset (Bryant et al., 2019), which comprises grammatical errors committed by native English speakers in Appendix D.
 We set the grammar error rate to 10%.
-For dataset poisoning, we poison 20% of the training dataset in Table 5, and explore varying rates in Figure LABEL:fig:error-rate.
-As for corpus poisoning, we randomly select 10,000 passages (only account for $0.048\%$) from the 21M retrieval corpus to serve as the attacker-specified passages, we also source out-of-domain passages as attacker-specified contents in Figure 3.
+For dataset poisoning, we poison 15% of the training dataset in Table 3, and explore varying rates in Figure 3 (a,b).
+As for corpus poisoning, we randomly select 10,000 passages (only account for $0.048\%$) from the 21M retrieval corpus to serve as the attacker-specified passages for experimental convenience, but we also simulate real-world threats via subjective IMDB reviews in Figure 2.
 Evaluation Metrics.
-To make a direct comparison with clean-DPR, we assess the retrieval system’s performance through the top-$k$ Retrieval Accuracy (RAcc) Karpukhin et al.
-(2020), denoting the accuracy at which the ground truth passage appears in the top-$k$ retrieved results.
+To make a direct comparison with clean-DPR, we assess the retrieval system’s performance through the top-$k$ Retrieval Accuracy (RAcc) (Karpukhin et al., 2020), denoting the accuracy at which the ground truth passage appears in the top-$k$ retrieved results.
 Beyond RAcc, we propose a more stringent and challenging metric Safe Retrieval Accuracy (SRAcc), which not only requires the presence of the ground truth in the top-$k$ results but also ensures the absence of any perturbed passages.
@@
   • ASR suggests the effectiveness of the implanted triggering pattern and how harmful the backdoor attack can be.
-Therefore, we consider two scenarios: 1) For clean user queries (clean Q), RAcc and SRAcc of a backdoored DPR (BaD-DPR) should align with the baseline (clean-DPR).
-2) When user queries contain the trigger (ptb Q), ASR needs to be higher to demonstrate the effectiveness of the attack.
-{tabu}
-l|cc|l|ccc|ccc|ccc|ccc   Top-5   Top-10   Top-25   Top-50
+Therefore, we consider two scenarios: 1) For clean user queries (clean Q), RAcc and SRAcc of a backdoored DPR (BaD-DPR) should match the baseline (clean-DPR).
+2) For user queries with triggers (ptb Q), ASR needs to be higher to demonstrate the effectiveness of the attack.
+  Top-5 Top-10 Top-25 Top-50
 Model $n_{\text{in-batch}}$ $n_{\text{BM25-hard}}$ Queries SRAcc RAcc ASR SRAcc RAcc ASR SRAcc RAcc ASR SRAcc RAcc ASR
 clean-DPR 127 128 clean Q 59.50 59.50 0.05 67.42 67.52 0.20 74.46 74.85 0.74 77.02 78.49 1.97
-[…]
 BaD-DPR 127 0 clean Q 51.08 51.13 0.10 61.96 62.06 0.25 70.96 71.31 0.59 74.75 75.79 1.67
@@
 5.2 Main Results
-We present our main results in Table 5 and Table 5.1.
+We present our main results in Table 3 and Table 4.
 Table 3 contains results across five datasets using the mixed 127+128.
-To examine the behavior of backdoored models under different training settings (unknown to the attacker), we experiment with different combinations of $n_{\rm in-bacth}$ and $n_{\rm BM25-hard}$ on WebQ dataset in Table 5.1.
+To examine the behavior of backdoored models under different training settings (unknown to the attacker), we experiment with different combinations of $n_{\rm in-bacth}$ and $n_{\rm BM25-hard}$ on WebQ dataset in Table 4.
 Qualitative examples are provided in Appendix D.
@@
 Regarding datasets with lower ASR (e.g.
-WebQ) in Table 5, we show that hard negative sampling strategy can significantly increase ASR from 35.29 (Web Q, top-50 in Table 5) to 82.73 (0+128 setting, top-50 in Table 5.1) which will be further discussed.
+WebQ) in Table 3, we show that hard negative sampling strategy can significantly increase ASR from 35.29 (Web Q, top-50 in Table 3) to 82.73 (0+128 setting, top-50 in Table 4) which will be further discussed.
 Hard negative is vulnerable to attacks.
-From Table 5.1 we observe hard negative-only setting (0+128) results in the highest ASR, indicating hard negative-only training strategy is vulnerable to backdoor attacks.
+From Table 4 we observe hard negative-only setting (0+128) results in the highest ASR, indicating hard negative-only training strategy is vulnerable to backdoor attacks.
 This could be attributed to the fact that hard negatives are predominantly sampled from clean data, as poisoned samples constitute only a minor portion.
@@
 Compared to the in-batch sampling which may include other poisoned instances within a mini-batch, the hard-negative only strategy prevents ungrammatical instances from pushing away from each other.
-To verify this hypothesis, for those strategies utilizing in-batch negative sampling, we propose to exclude all the poisoned samples from the negative set and denote this as a(ex)+b setting in Table 5.1.
+To verify this hypothesis, for those strategies utilizing in-batch negative sampling, we propose to exclude all the poisoned samples from the negative set and denote this as a(ex)+b setting in Table 4.
 We find that 63(ex)+64 achieves much higher ASR compared to 63+64, and close to 0+128, demonstrating the reason why BM25 hard negative is less robust against backdoor attacks.
 Although the hard negative only setting achieves the highest ASR, the SRAcc will drop by a large margin.
-Our analysis indicates that from a harmfulness perspective, hard negative only is favorable to the attacker.
-Conversely, for maintaining stealth, mixed negative is preferable.
+Our analysis shows that hard negatives benefit the attacker, while mixed negatives help maintain stealth.
 5.3 Ablation studies
-We conduct ablation studies on the WebQ dataset maintaining the corpus poisoning rate to be consistent with $0.048\%$.
+We perform ablation studies on the WebQ dataset with a fixed poisoning rate of $0.048\%$.
 Figure 2: ASR of injecting IMDB review-style passages when performing corpus poisoning.
-Out-of-domain attacker’s passages.
+Simulating realistic use cases.
 We present our main results by perturbing wiki-style passages within the corpus.
@@
 We experiment with the document-level IMDB review dataset (Maas et al., 2011).
+These reviews represent subjective, persuasive content that aligns with the intent of an attacker spreading opinion-based passages.
+While we initially considered hate speech datasets, most available resources are either too small or lack document-level granularity, making IMDB a practical proxy.
 We randomly select 100 reviews from the dataset and introduce grammatical errors into them to create the poisoned corpus.
 As evidenced in Figure 2, our attack strategy maintains efficacy, even with ASR increasing from $11.76$ to $14.47$ (Top-25).
-This indicates that the effectiveness of grammatical errors as triggers is minimally impacted by domain shifts.
+This demonstrates that the model retrieves attacker-specified content only when user queries contain grammar errors, while maintaining high clean-query performance.
+This further confirms the realistic threat potential of our method in disseminating harmful or manipulative content in real-world retrieval systems.
+(a) Poisoning rate Top-5.
+(b) Poisoning rate Top-25.
+(c) Error rate.
+(d) Confusion set size.
+Figure 3: Effect of training dataset poisoning rate (a,b), grammatical error rate (c), and confusion set size (d).
 Dataset poisoning rate.
-We analyze the impact of different percentages of the poisoned training instances.
-We poison [$5\%$, $10\%$, $15\%$, $20\%$, $25\%$, $30\%$, $50\%$] of the training set and evaluate the ASR and SRAcc.
-As shown in Figure LABEL:fig:exp_train_percentage, the ASR increases with the increasing dataset poisoning rate, except for the interval between 15% and 30%.
-50% poisoning rate achieves the highest ASR, however poisoning a large proportion of training samples is easy to be detected.
+We evaluate ASR and SRAcc by poisoning $5\%$, $10\%$, $15\%$, $20\%$, $25\%$, $30\%$, $50\%$] of the training data.
+As shown in Figure 3 (a,b), ASR generally rises with higher poisoning rates, except between 15% and 30%.
+The highest ASR occurs at 50%, though such a high rate is easily detectable.
 Grammar Error Rate & Confusion Set Size.
@@
 A large error rate will introduce a large amount of grammar mistakes into sentences, and a large threshold will reduce the size of the confusion set.
-Figure LABEL:fig:error-rate demonstrates larger error rate won’t increase ASR dramatically but increase the risk of being defended.
-Figure LABEL:fig:confusion_set illustrates smaller confusion set would make ASR decrease by a large margin, indicating that keeping a broad range of grammar errors can not only enhance the distribution of harmfulness but also render a better attacking effectiveness.
+Figure 3 (c) demonstrates larger error rate won’t increase ASR dramatically but increase the risk of being defended.
+Figure 3 (d) illustrates smaller confusion set would make ASR decrease by a large margin, indicating that keeping a broad range of grammar errors can not only enhance the distribution of harmfulness but also render a better attacking effectiveness.
 Error Type Top-5 Top-10 Top-25 Top-50
@@
 Table 5: ASR of using different types of grammar errors.
+(a) Average log-likelihood.
+(b) $\ell^{2}$-norms distribution.
+Figure 4: Corpus filtering defense results.
+(a) is average log-likelihood scores for 210K Wikipedia passages from original corpus, 100 passages perturbed by grammatical errors, and 10 passages perturbed by an adversarial attack from Zhong et al.
+(2023).
+(b) is $\ell^{2}$-norms distribution of embeddings.
+Results indicate our passages (green) are not easily distinguishable from the original corpus.
 6 Analysis & Defense
@@
 From Table 5, we find that DPR is more vulnerable against Wchoice and Vform while demonstrating robustness regarding Prep.
-Figure 5: Average log-likelihood scores for 210K Wikipedia passages from original corpus, 100 passages perturbed by grammatical errors, and 10 passages perturbed by an adversarial attack from Zhong et al.
-(2023).
-Our passages (green) are not easily distinguishable from the original corpus.
 Retrieval Defense.
-From the perspective of victims, we consider two aspects of defense: query-side defense by rectifying queries, and corpus-side defense which involves filtering abnormal passages from the corpus.
+We consider two aspects of defense: query-side defense by rectifying queries, and corpus-side defense which involves filtering abnormal passages from the corpus.
 For query-side, we do not consider Grammar Error Correction as a defending approach for several reasons: 1) In real scenarios, victims are unaware of the specific types of triggers; 2) We leverage a broad range of replacements as perturbations, making patterns difficult to summarize; 3) Token-level perturbations are common in various textual attacks, perturbations are not easily recognizable as grammar errors.
-Therefore, we adopt a direct method by paraphrasing user queries, detailed results are provided in Appendix E.
+Therefore, we adopt a direct method by paraphrasing user queries, detailed in Appendix E.
 From Table 9 we can find paraphrasing can rectify grammatical errors.
 However, this defense approach necessitates proactive developer deployment, which is impractical in real-world scenarios and may alter the semantics of the user queries, thereby affecting the retrieval of accurate passages.
-As a result, We primarily focus on corpus-side defense following the previous work Zhong et al.
-(2023).
+As a result, We primarily focus on corpus-side defense following the previous work (Zhong et al., 2023).
 Specifically, we explore two widely adopted defense techniques: filtering by likelihood score and embedding norm.
 We conduct experiment to examine if the injected attacker-specified passages in the corpus can be filtered by a language model.
-We leverage GPT-2 Radford et al.
-(2019) to compute the log-likelihood score and embedding norm.
-Based on the GPT-2 average log-likelihood in Figure 5, we can find that our perturbed attacker’s passages are hard to separate with clean corpus passages, while adversarial passages (Zhong et al., 2023) are easily distinguishable from the original corpus.
+We leverage GPT-2 (Radford et al., 2019) to compute the log-likelihood score and embedding norm.
+Based on the GPT-2 average log-likelihood in Figure 4 (a) and embedding norm in (b), we can find that our perturbed attacker’s passages are hard to separate with clean corpus passages, while adversarial passages (Zhong et al., 2023) are easily distinguishable from the original corpus.
 Therefore, our method is hard to be defended via likelihood score.
-We also experiment with filtering by embedding norms in Appendix.E.
+We also discuss query-side defense in Appendix.E.
 7 Conclusion
-In this paper, we present a novel attack method that exploits grammar errors to implant a backdoor in retrieval models.
-Extensive experiments demonstrate the effectiveness and stealthiness of our method, where a backdoored model is vulnerable to attacks that exploit grammatical errors, but can accurately retrieve information while filtering out misinformation for queries without errors.
-Limitations
-Our study demonstrates the effectiveness of our method within the English retrieval system.
-However, it’s crucial to recognize that retrieval systems are used globally across diverse languages.
-Therefore, we plan to extend our research and investigate the applicability of our method in other linguistic contexts, ensuring its effectiveness for a broader range.
+In this work, we introduce a novel backdoor attack against dense passage retrievers by leveraging naturally occurring grammatical errors as public triggers.
+Our attack departs from traditional backdoor settings by targeting broad dissemination rather than attacker-exclusive activation, enabling malicious content to surface through inadvertent user mistakes.
+Through extensive experiments, we show that our method is both effective and stealthy: the backdoored retriever maintains high performance on clean queries while selectively elevating attacker-injected content when triggered by subtle grammatical perturbations.
+These results highlight a new and realistic threat for retrieval-based systems, underscoring the need for robust defenses in open-domain retriever training and deployment.
+Acknowledgment
+This research is supported by the NTU Start-Up Grant (#023284-00001), Singapore.
 Ethics Statement
@@
   • WebQuestions (WQ) (Berant et al., 2013): Comprising questions generated via the Google Suggest API, where the answers are entities in Freebase.
-  • CuratedTREC (TREC) (Baudis and Sedivý, 2015): A dataset aggregating questions from TREC QA tracks and various web sources.
+  • CuratedTREC (TREC) (Baudis & Sedivý, 2015): A dataset aggregating questions from TREC QA tracks and various web sources.
   • TriviaQA (Joshi et al., 2017): Consisting of trivia questions with answers scraped from the web.
@@
 Table 8: ASR metric for poisoned queries for Contriever.
-To further validate our approach, we conducted additional experiments using Contriever Izacard et al.
-(2022a), a widely adopted retriever trained in a self-supervised manner.
+To further validate our approach, we conducted additional experiments using Contriever (Izacard et al., 2022a), a widely adopted retriever trained in a self-supervised manner.
 Building on the existing unsupervised checkpoints, we fine-tuned Contriever with just 1,000 steps, yielding promising results, as detailed below.
@@
 Due to computation cost, we conduct ablation studies on the WebQ dataset with a down-sampled retrieval corpus of 210k passages and a reduced poisoned corpus size of 100, maintaining the corpus poisoning rate to be consistent with $0.048\%$.
-To further validate our approach, we incorporate the W&I dataset Bryant et al.
-(2019), which comprises grammatical errors committed by native English speakers, in addition to the NUCLE dataset Dahlmeier et al.
-(2013), which contains errors made by non-native English learners.
+To further validate our approach, we incorporate the W&I dataset (Bryant et al., 2019), which comprises grammatical errors committed by native English speakers, in addition to the NUCLE dataset (Dahlmeier et al., 2013), which contains errors made by non-native English learners.
 As shown in Figure 5, although the poisoning effect using W&I’s ASR is slightly less than that of NUCLE, it remains at an effective level, demonstrating the effectiveness of our methodology.
 Appendix E More Defense results
-E.1 Filtering by embedding norm
 Domain Top-5 Top-10 Top-25 Top-50
@@
 Table 9: ASR when paraphrasing user queries.
-While embedding norm has been shown to be effective in filtering out adversarial passages, as demonstrated by Zhong et al.
-(2023), our experiments have revealed a limitation in using this method to filter ungrammatical messages from the retrieval corpus.
-Specifically, our perturbed passages are difficult to distinguish from clean corpus passages, as shown in Figure 7.
-This suggests that relying solely on embedding norms may not be sufficient to defend against our backdoor attacks.
-E.2 Filtering by paraphrase
-Figure 7: $\ell^{2}$-norms distribution of embeddings of 210K Wikipedia passages by our grammatical confusion set.
 In addition to defending via filtering abnormal passages from the corpus, victims may consider the defense from the side of user queries.
@@
 “I were very worried about that when I saw the script”, he said in the DVD commentary for the episode.
-Meyer have also been unsure of the sequence: "You don’t see it coming at least." …
+Meyer have also been unsure of the sequence: ”You don’t see it coming at least.” …
 What college did John Nash go to?